### v2

# Hierarchical Vision-Language Prompting for Long-form Video Understanding

Abstract
Long-form videos that span across wide temporal intervals are highly information-redundant and contain multiple distinct events or entities that are often loosely-related. Therefore, when performing long-form video question answering (LVQA), all information necessary to generate a correct response can often be contained within a small subset of frames. Recent literature explore the use of large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language. Such VLMs often independently caption a large number of frames uniformly sampled from long videos, which is not efficient and can mostly be redundant. Questioning these decision choices, we explore optimal strategies for key-frame selection and sequence-aware captioning, that can significantly reduce these redundancies. We propose two novel approaches that improve each of aspects, namely Hierarchical Keyframe Selector and Sequential Visual LLM. Our resulting framework termed LVNet achieves state-of-the-art performance across three benchmark LVQA datasets. Our code will be released publicly.
1 Introduction
Video understanding is a long-standing vision problem [Aggarwal2011HumanAA] with numerous real-world applications. It has been traditionally studied even before the era of differentiable representation learning, with hierarchical approaches focusing on longer videos [allen1994actions, ivanov2000recognition, shi2004propagation, hongeng2004video, Ryoo2006RecognitionOC]. Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering that involve generating human-style conversations in response to questions regarding videos [dataset_tapaswi2016movieqa, videoqaaaai2017, xu2017msvdqavideo].
The recent popularity of vision-language models (VLMs), particularly approaches connecting large language models (LLMs) to vision architectures [liu2023llava, li2023blip, dai2023instructblip], has resulted in significant improvements across visual question answering (VQA) tasks. These models demonstrate exceptional performance within the image domain, and their video variants [Yu2023SelfChainedIM, papalampidi2023simple, Maaz2023VideoChatGPT] perform similarly on shorter videos, yet demonstrating limited performance on long-form video benchmarks [Mangalam2023EgoSchemaAD, Kahatapitiya2024langrepo, rana2024mvu, rawal2024cinepile]. This can be attributed to the nature of long-form video benchmarks, which require both temporal sequence awareness and causal reasoning. An alternate line of works [zhang2023llovi, wang2023vamos, Kahatapitiya2024langrepo, Wang2024VideoAgentLV] adapt LLMs that contain strong reasoning abilities for this task, using image VLMs to generate per-frame natural language descriptions, followed by video question answering purely within the language domain. However, these methods employ expensive VLMs to caption a large number of uniformly sampled frames. Such a design choice leading to high compute expense, is questioned in [Buch2022RevisitingT, rana2024mvu, Wang2024VideoAgentLV], and is the key motivation for our exploration of key frame selection, i.e. identifying a minimal set of frames most useful for correctly answering a given video-question pair. Moreover, existing captioning methods treats each frame independently while overlooking their sequential nature, which can result in significant redundancies and lack of compatibility between generated captions. This leads to our secondary motivation of exploring sequence-aware captioning strategies for videos.
Figure 1: (left) Our proposed LVNet contains a novel Hierarchical Keyframe Selector module and Sequential Visual LLM strategy followed by an LLM for answer generation. (right) LVNet achieves competitive performance using a fraction of captioned frames highlighting the value of optimal key frame selection.
Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) and Sequential Visual LLM (SVL) that performs efficient key-frame selection followed by sequence-aware captioning of individual frames as illustrated in Figure 1. Aligned with prior work [zhang2023llovi, wang2024lifelongmemory, Wang2024VideoAgentLV], the per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair. The scope of this work focuses on optimizing the prior two stages.
We summarize our key contributions as follows:
  • Long-form Video Understanding framework requiring zero video-level training,
  • Novel Hierarchical Keyframe Selector module,
  • Sequence-aware frame captioning strategy (Sequential Visual LLM module).
Our proposed LVNet achieves state-of-the-art results on three long-form video question answering benchmarks, EgoSchema [Mangalam2023EgoSchemaAD], Next-QA [xiao2021nextqa], and IntentQA [li2023intentqa_cavir], demonstrating strong performance and general applicability of our approach.
2 Related Work
Video Question Answering: Visual question answering (VQA), a relatively recent task, involves generating free-form and open-ended textual content conditioned on an image and natural language text [Agrawal2015VQAVQ]. Its video variant, Video-VQA [Yu2019ActivityNetQAAD] replaces images with videos. While multiple early datasets focus on querying objects or events based on referential and spatial relations [xu2017msvdqavideo, videoqaaaai2017, Yu2019ActivityNetQAAD], later tasks require explicit temporal understanding of sequential events [dataset_lei2018tvqa, lei-etal-2020-tvqaplus, dataset_anetqa]. More recent datasets focus on longer videos containing multiple actions and scenes spread over a wide time interval (termed long-form videos) [dataset_xiao2021nextqa, dataset_li2022causalvidqa] with questions additionally requiring causal reasoning to correctly answer. Referred to as long-form video question answering (LVQA), these benchmarks are constructed to specifically test strong causal and temporal reasoning [dataset_xiao2021nextqa] over long temporal windows [Mangalam2023EgoSchemaAD]. Some works tackling such casual video VQA tasks leverage graph networks to model cross object / event relations [lmappce1_hosseini-etal-2022-knowledge, model_xiao2021hgqa, model_xiao2022vgt]. A more recent line of works integrate large-language models (LLMs) to tackle this task [zhang2023llovi, wang2023vamos, Kahatapitiya2024langrepo, Wang2024VideoAgentLV, rana2024mvu, wang2024lifelongmemory] utilizing the strong reasoning skills of the LLMs. A common aspect is the use of a vision langauge model (VLM) to convert the frame level visual information into natural language. This is in turn input to the LLM which makes a final prediction. In contrast to these methods, our LVNet contains a) a unique Hierarchical Keyframe Selector and b) a novel Sequential Visual LLM strategy for frame captioning. We also highlight that unlike early supervised methods, our LVNet requires no video level training.
Frame Selection in Videos: The task of frame selection in videos has been long explored in video [davis1997mei, Zhao_2017_ICCV] with more recent works focused directly on long-form video question answering [Buch2022RevisitingT, rana2024mvu]. Most similar to our work is [Wang2024VideoAgentLV] which employs an LLM based strategy for video frame selection. However, our proposed Hierarchical Keyframe Selector module differs with unique visual only operations as well as cross-modal operations handling much longer frame sequences extending up to 900 frames.
3 Method
In this section, we present our training-free (i.e. zero-shot) framework for long-form video QA, LVNet. Videos are a dense form of data with even a few seconds long clip being composed of 100s of frames (individual images). In the case of long-form videos, this frame count is even greater. However, the information necessary to answer a given question is often contained in a handful of those frames. Our framework tackles this challenge of selecting an optimal and minimal set of informative frames. We refer to this as keyframe selection. Given such a set of useful frames, we also establish optimal strategies for extracting their information using modern large language models (LLMs), taking into account their sequential nature.
Our proposed LVNet comprises of three components: a Hierarchical Keyframe Selector (HKS), a Sequential Visual LLM (SVL), and an LLM as illustrated in Figure 1. The proposed HKS, an efficient, hierarchical keyframe selector is our key contribution. HKS processes up to 900 frames per video conditioned on a given natural language query in order to select a subset of frames most useful for answering that question. The selected frames are described in natural language using a more expensive, powerful vision language mode (VLM). Our second contribution is a strategy that accounts for the sequential nature of these frames when captioning them using a VLM. Finally, an LLM processes all language descriptions of the selected frames to answer a given query.
3.1 Background
Recent approaches utilizing LLMs for long video question answering (LVQA) [zhang2023llovi, wang2023vamos, Kahatapitiya2024langrepo, rana2024mvu, Wang2024VideoAgentLV] can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation. Note that the complexity of each stage varies across methods given their focus on different aspects of the LVQA task (e.g. frame selection in some is simply uniform sampling). In our work, we also follow this structure, but we focus on improving the first two stages with emphasis on frame selection.
Under such a framework, our proposed HKS and SVL modules can serve as plug-in modules to replace the frame selection and VLM based frame captioning stages respectively. Our final LLM based answer selection stage operates similar to these prior works.
3.2 Architecture
Consider a long video, $\mathbf{x}\in\mathbb{R}^{T\times C\times H\times W}$ with T, C, H, W for frames, channels, height, and width respectively. Its paired natural language query is referred as $\mathbf{q}$. Also consider a frame in $\mathbf{x}$ at timestamp $t$ as $\mathbf{x}[t]\in\mathbb{R}^{C\times H\times W}$. Our goal is to output the suitable response to this query based on information contained in the video. We refer this correct response as $\mathbf{r}$.
Our LVNet processes a given video-query ($\mathbf{x},\mathbf{q}$) pair to output a response, $\mathbf{\hat{r}}$. The HKS module first processes this video-query pair, selects $T^{\prime}$ keyframes, and outputs a deterministically sub-sampled video $\mathbf{x^{\prime}}\in\mathbb{R}^{T^{\prime}\times C\times H\times W}$. Each of these $T^{\prime}$ frames are then processed by our SVL stage to generate a set of natural language descriptions, $D=\{d_{1},d_{2},...d_{T^{\prime}}\}$ where $d_{i}$ describes the frame $\mathbf{x^{\prime}}[i]$. Finally, the LLM processes all descriptions $D$ and the query $\mathbf{q}$ to generate response $\mathbf{\hat{r}}$.
Our unique contributions here are the Hierarchical Keyframe Selector (HKS) and the Sequential Visual LLM (SVL). We also identify optimal prompts and description templates for the LLM when generating the final response which is also important as suggested by our experiments. The overall architecture is given in Figure 1 (left).
3.3 Hierarchical Keyframe Selector
Figure 2: Detailed architecture of Hierarchical Keyframe Selector: Input frames first go through a Temporal Scene Clustering (TSC) mechanism, which divides the long-video into scenes, enabling per-scene subsampling. Next, a Coarse Keyframe Detector (CKD) is applied to select frames best-aligned with keywords relevant to the query. Finally, a Fine Keyframe detector further subsamples frames, by refining keyword alignements through a templated visual prompting.
We now describe our proposed Hierarchical Keyframe Selector (HKS) module. As illustrated in Figure 2, our HKS comprises of three submodules, namely 1) Temporal Scene Clustering (TSC), 2) Coarse Keyframe Detector (CKD), and 3) Fine Keyframe Detector (FKD). The inputs to HKS contain $T$ frames with each of the three stages TSC, CKD, and FKD reducing this frame count to $T_{a}$, $T_{b}$, and $T_{c}=T^{\prime}$ respectively.
Temporal Scene Clustering (TSC): The role of TSC is to perform visual content aware preliminary frame sampling. The established approach for preliminary frame selection is uniform sampling. In contrast, TSC extracts per-frame visual features using a deep neural network followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets. Within each of the n sets, we uniformly sample $k$ frames obtaining a total of $T_{a}=k\times n$. Our iterative clustering procedure is outlined in LABEL:alg:tsc. It calculates pairwise distances between all frames accounting for intra-frame local information using the extracted per-frame features, followed by n iterative frame similarity based grouping operations. Further details in LABEL:app:icp.
Our reasoning for TSC is that videos, especially long-form ones, contain extensive information redundancy across its temporal dimension. For example, in a long sub-sequence of mostly static frames from a single scene, beyond its first frame the rest may contain little additional information. We hypothesize that our TSC module will utilize frame level feature similarity to cluster the frames, such that each cluster corresponds to a common scene or motion. A single cluster could contain just one frame or significantly more based on frame feature similarities. This leads to a non-uniform sampling of frames across the entire video. In contrast to uniform sampling across the entire videos where each timestamp is treated equally, in our formulation the uniform sampling within segments will allow more frames to be sampled from information heavy temporal regions.
Coarse Keyframe Detector (CKD): Unlike the TSC in prior stage, CKD reasons across both visual and language modalities (using the paired textual query) to further sub-sample the $T_{a}$ frames output from the TSC module. The inputs to this CKD stage are the $T_{a}$ frames and the paired query $\mathbf{q}$. The module generates a set of keywords $K_{\text{CKD}}$, selects $T_{b}$ frames conditioned on these keywords, and outputs the selected $T_{b}$ frames with a confidence value and keyword assigned to each frame.
The CKD contains three elements: a keyword generation strategy, a dual-encoder image-text model, and an iterative algorithm for similarity based confidence assignment. The keyword generation strategy utilizes the given query, q alongside a hand-crafted templating operation or an LLM to select or generate suitable keywords. The dual-encoder image-text model uses a spatially aware contrastive langauge image pre-training (CLIP) network from [Ranasinghe2022PerceptualGI]. For confidence assignment, we construct an algorithm as outlined in LABEL:alg:kwimg which processes two lists, one of frames and one of keywords, and then calculates their pairwise likelihood of occurrence to assign each frame a confidence value (that reflects its usefulness to answer the query, $\mathbf{q}$).
The keyword generation strategy converts the query $\mathbf{q}$ into a set of keywords, $K_{\text{CKD}}$ where the number of keywords $|K_{\text{CKD}}|\in[8,25]$. The dual-encoder model converts the keyword set $K_{\text{CKD}}$ and the frame set (containing $T_{a}$ frames) into two sets of embeddings. The embeddings of each set are comparable to each other using a cosine similarity distance metric given the training formulation of the CLIP model. These two embedding sets are processed by our confidence assignment algorithm in LABEL:alg:kwimg which assigns each of the $T_{a}$ frames a confidence score and a keyword. Note that the same keyword can be assigned to multiple frames. However, given the much higher frame count compared to the keyword count and the cross-frame redundancy, we do not assign multiple keywords to a single frame.
The key role of this module is language conditioned frame selection. For a single query, there can be multiple regions in a video that are highly informative but not useful or relevant in answering that query. A single query can also contain multiple different concepts and attributes that must be given attention to construct a correct answer: the keyword generation attempts to capture each of these distinct attributes. On the visual modality, a single frame will also encode multiple concepts and attributes. Our design choice for the spatially aware CLIPpy dual-encoder VLM from [Ranasinghe2022PerceptualGI] is motivated by this nature of individual frames. Finally, confidence assignment takes into account these multiple modes of information within each frame and the query and suitably assigns a confidence score to each frame that reflects its relevance to the query. We also highlight how the confidence scores are directly linked to the related keyword (i.e. reason that makes the frame relevant) leading to better interpretability and the ability to perform further keyword based refinement in later stages.
Fine Keyframe Detector (FKD): In the prior CKD stage, cross-modal selection utilizes a dual-encoder VLM that is constrained by the set of keywords provided and performs limited reasoning at frame level. In contrast, FKD comprises of a) a visual templating module to combines multiple frames and b) an autoregressive VLM that can generate open-ended natural language outputs and is capable of more higher level reasoning. The input to this stage is the set of $F_{b}$ frames, with each frame having an assigned confidence score and keyword.
Our visual templating module partitions the $T_{b}$ frames into sets of 8 ordered by their confidence scores, arranges frame sets as grids to form a collage-style image, and annotates that image with visually identifiable tags corresponding to each frame. We further illustrate this process in Figure 3 (see Visual Templating column). Each of these visual templated images also contain a subset of the keywords that correspond to their 8 images. These resulting visual templated images along with a prompt containing their associated keywords and instructions to select a frame subset based on valid association between keywords and images (see LABEL:app:prompt_fkd for details) are input to the VLM. The output of the VLM is used to select a subset of each 8 image group. These frames are collected as the output of the FKD stage, overall resulting in $T_{c}$ frames.
The purpose of the initial visual templating module is to allow reasoning across a set of frames using the image-text VLM (which is trained to process a single image at at time). This partitioning of the input $T_{b}$ frames is performed based on confidence scores from the prior stage. The eight frames with top confidence scores are grouped into the first visual template, followed by the next eight and so forth. This ensures the VLM selects both high confidence concepts and low confidence concepts, accounting for biases and weaknesses in our CKD stage VLM. Our intuition is that such a mechanism allows to best utilize the complimentary strengths of two different VLMs from CKD and FKD stages for better frame selection overall.
3.4 Sequential Visual LLM
Our secondary contribution is a strategy for frame caption generation taking into account their sequential nature. We refer to this module as Sequential Visual LLM (SVL).
In contract to the standard approach of independent frame captioning, we explore how per-frame caption generation can be conditioned on captions of previous frames. The VLM used to generate frame descriptions is additionally provided with the caption of one previous frame alongside suitable prompting and templating (details in LABEL:app:vlm).
This is motivated by accounting frame sequential order as well as the presence of redundancy across frames. We discover how this minimal modification to the caption generation strategy results in improved captions, both avoiding caption redundancy and focusing on subtle details in videos (see examples in LABEL:app:vlm).
4 Experiments
In this section, we first discuss our experimental setup followed by quantitative evaluations comparing to existing baselines and ablations of our proposed components. We then present qualitative results for our method and outline some limitations of our approach.
4.1 Experimental Setup
Datasets: Given the training free nature of our framework, we do not utilize any video datasets for training. Datasets are used purely for evaluation. We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema [Mangalam2023EgoSchemaAD], NExT-QA [dataset_xiao2021nextqa], and IntentQA [li2023intentqa_cavir]. EgoSchema is a very long-form video question-answering dataset derived from Ego4D. For each question, EgoSchema requires the correct answer to be selected between five given options based on a three-minute-long video clip. We evaluate on both its 500 subset and the 5000 fullset.
NExT-QA is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer. We evaluate on its validation set containing 4996 questions over 570 videos. The third dataset, IntentQA [li2023intentqa_cavir] is based on NExT-QA videos corresponding to temporal and causal reasoning quetions. It consists of 16k multiple-choice questions based on 4.3k videos, each with 5 answer choices. These questions are further classified as why?, how? or before/after, based on the appearance of the same keywords in the question.
VLM and LLM Model Choices: In line with prior state-of-the-art work [wang2024lifelongmemory, zhang2023llovi, wang2023vamos] we utilize GPT-4 API for the VLM and LLM in the latter two stages of our framework. This allows fairer comparison against these methods. In the case of the HKS module, we use the same VLM / LLM for keywords used in CKD and VLM used in FKD.
4.2 Quantitative and Qualitative Evaluation
Table 1: Results on EgoSchema [Mangalam2023EgoSchemaAD]: We present comparisons with state-of-the-art models on EgoSchema subset (500-videos) and fullset (5000-videos). LVNet shows a strong zero-shot performance using only a fraction of frames compared to most methods.
Model Captions Subset (%) Fullset (%)
finetuned
ImageViT [papalampidi2023simple] 16 40.8 30.9
ShortViViT [papalampidi2023simple] 32 47.9 31.0
MC-ViT-L [balavzevic2024mcvit]   128+ 62.6 44.4
LongViViT [papalampidi2023simple] 256 56.8 33.3
zero-shot
VIOLET[fu2023violet] 5 - 19.9
mPLUG-Owl [ye2023mplug] 5 - 31.1
VideoAgent [Wang2024VideoAgentLV] 8.4 60.2 54.1
MVU [rana2024mvu] 16 60.3 37.6
SeViLA [yu2024sevila] 32 25.7 22.7
LifelongMemory [wang2024lifelongmemory] 39 26.4 -
InternVideo [wang2022internvideo] 90 - 32.1
FrozenBiLM [yang2022frozenblim] 90 - 26.9
LifelongMemory [wang2024lifelongmemory] 90 68.0 62.1
LLoVi [zhang2023llovi] 180 57.6 50.3
LangRepo [Kahatapitiya2024langrepo] 180 66.2 41.2
Vamos [wang2023vamos] - - 48.3
LVNet (ours) 12 66.0 61.1
We first evaluate on the EgoSchema dataset and present our results in Table 1. Our framework outperforms most prior work to create a stronger baseline with a few frame captions. While LifelongMemory [wang2024lifelongmemory] achieves the absolute best zero-shot performance on the subset at 90 captions (68.0%), its performance drops significantly at 39 captions (26.4%). In contrast, LVNet shows a 66.0% accuracy, merely at 12 captions, even outperforming VideoAgent [Wang2024VideoAgentLV] (60.2% at 8.4 captions) and MVU [rana2024mvu] (60.3% at 16 captions). We achieve 61.1% for the fullest, the highest among the models utilizing approximately 12 frames. Here, we generate captions independently without conditioning on previous captions and present them to the LLM in a system-user-prompt style.
Table 2: Results on NExT-QA [xiao2021nextqa]: We compare LVNet against prior methods on NExT-QA validation set, highlighting standard splits: causal, temporal and descriptive. Our method shows state-of-the-art zero-shot performance, even competing with finetuned models.
Model Captions  Causal (%)  Temporal (%)  Descriptive (%)  All (%)
finetuned
HiTeA [ye2023hitea] 8 62.4 58.3 75.6 63.1
SeViTFiD [kim2023sevit] 10 - - - 60.6
BLIP-2 [li2023blip2] 10 70.1 65.2 80.1 70.1
LLama-VQA [ko2023llama-vqa] 10 72.7 69.2 75.8 72.0
CoVGT [xiao2023covgt] 32 58.8 57.4 69.3 60.0
SeViLA [yu2024sevila] 32 74.2 69.4 81.3 73.8
InternVideo [wang2022internvideo] 90 62.5 58.5 75.8 63.2
MC-ViT-L [balavzevic2024mcvit] 128+ - - - 65.0
Vamos [wang2023vamos] - 72.6 69.6 78.0 72.5
zero-shot
VideoAgent [Wang2024VideoAgentLV] 8.2 72.7 64.5 81.1 71.3
MVU [rana2024mvu] 16 55.4 48.1 64.1 55.2
VFC [momeni2023vfc] 32 45.4 51.6 64.1 51.5
SeViLA [yu2024sevila] 32 61.3 61.5 75.6 63.6
InternVideo [wang2022internvideo] 90 43.4 48.0 65.1 49.1
LLoVi [zhang2023llovi] 90 69.5 61.0 75.6 67.7
LangRepo [Kahatapitiya2024langrepo] 90 64.4 51.4 69.1 60.9
ViperGPT [suris2023vipergpt] - - - - 60.0
LVNet (ours) 12 75.0 65.5 81.5 72.9
We next evaluate on the Next-QA dataset and present results in Table 2. This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs. Our approach achives state-of-the-art performance on this benchmark outperforming prior work by a considerable margin. In fact, our LVNet with just 12 frame captions achieves 72.9% overall accuracy in zero-shot, outperforming VideoAgent [Wang2024VideoAgentLV] (71.3% at 8.2 captions) and LLoVi [zhang2023llovi] (67.7% at 90 captions). It is also extremely competitive with finetuned models such as LLama-VQA [ko2023llama-vqa] (72.0% at 10 captions), SeViLA [yu2024sevila] (73.8% at 32 captions) and Vamos [wang2023vamos] (72.5%), while having no video level training.
Given the generative nature of VQA tasks as well as the limited availability and noisy nature of fully-annotated video VQA corpora, building generalizable fully-supervised models are challenging for these tasks. Nevertheless, we highlight how our zero-shot and video level training-free framework is competitive with the best supervised approaches on this dataset. This indicates the promise of utilizing pretrained models, especially those equipped with extensive world knowledge and reasoning skills from alternate modality specific learning (i.e. in our cases image domain VLMs and language domain LLMs).
Table 3: Results on IntentQA [li2023intentqa_cavir]: We compare LVNet against prior methods on IntentQA test set, highlighting standard splits: why?, how? and before/after. Our method shows state-of-the-art performance in zero-shot, with a significant margin.
Model Captions Why? (%) How? (%) Before/After (%) All (%)
finetuned
HQGA [xiao2022hqga] 16 48.2 54.3 41.7 47.7
VGT [xiao2022vgt] 32 51.4 56.0 47.6 51.3
BlindGPT [ouyang2022blindgpt] - 52.2 61.3 43.4 51.6
CaVIR [li2023intentqa_cavir] - 58.4 65.5 50.5 57.6
Vamos [wang2023vamos] - 69.5 70.2 65.0 68.5
zero-shot
SeViLA [yu2024sevila] 32 - - - 60.9
LLoVi [zhang2023llovi] 90 68.4 67.4 51.1 64.0
LangRepo [Kahatapitiya2024langrepo] 90 62.8 62.4 47.8 59.1
LVNet (ours) 12 75.2 71.6 60.8 71.1
We finally evaluate on IntentQA [li2023intentqa_cavir] dataset and report these results in Table 3. Our approach achieves state-of-the-art results on this benchmark as well outperforming all prior work, both zero-shot and finetuned. In fact, LVNet achieves a significant boost (+6.5%) over the closest zero-shot baseline LLoVi [zhang2023llovi] (64.0% at 90 captions), showing a 70.5% accuracy with just 12 frame captions. It also extends the state-of-the-art finetuned performance by +2.0%, compared to Vamos [wang2023vamos] (68.5%).
Figure 3: Qualitative example: We illustrate a challenging long-video QA scenario from EgoSchema [Mangalam2023EgoSchemaAD]. We consider an input of 900 frames, which first get clustered into scenes and subsampled to retain around 280 frames. Next, the Coarse Keyframe Detector selects only 32 frames out of them, based on the alignment with keywords (Here, keywords are extracted based on answer options, via an LLM). Such coarse keyframes are then ordered according to confidence values, and grouped into 4 groups of 8 frames each. These groups go through visal templating (i.e. simple concatenation across space) and fed into a VLM for Fine Keyframe Detection, resulting in just 12 frames.
4.3 Ablations
Table 4: Ablations on EgoSchema [Mangalam2023EgoSchemaAD]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot video VQA.
Keyframe sampler Seq. captions #captions Acc. (%)
Uniform Sampling (LLoVi) 180 57.6
HKS (w/o SVL) 12 63.0
HKS 12 63.4
Visual Templating order Acc. (%)
Temporal order 62.9
Confidence order 63.4
LLM Acc. (%)
GPT-3.5 59.2
GPT-4 61.1
GPT-4o 63.4
Frames 6.4 8.4 9 11 12 16
VideoAgent 58.4 60.2 - 57.4 - -
LVNet (ours) - - 64.6 - 66.0 65.8
In this section, we present ablations to key design decisions such as #frames, #words, prompting, and hierarchical processing. The results for these ablations are presented in Table 4. In all ablations, we consider a subset of 150 videos from EgoSchema [Mangalam2023EgoSchemaAD] for evaluation.
Keyframe sampler:
We evaluate our Hierarchical Keyframe Sampling strategy (HKS) against the established uniform sampling baseline from [zhang2023llovi]. Our results clearly indicate the importance of our HKS module.
Sequential captioning:
In our SVL module, we consider a sequential captioner, that considers previous captions when captioning the current frame. It outperforms the common parallel frame-wise captioning by $+0.4\%$ showing that the SVL can generate better captions, possibly reducing redundant information among captions and enabling a form of temporal modeling.
Visual Templating order:
In Visual Templating, grouping the frames based on confidence is better than preserving temporal order. It can allow sampling of challenging keyword frames, instead of only selecting high confidence keyframes.
Choice of LLM:
Here, we see that more-recent GPT-4o outperforms both GPT-4 and GPT-3.5. Since GPT-4o is also cheaper compared to GPT-4, we consider it as our default choice.
Number of frame captions:
We ablate different number of frame captions used in our setup compared to VideoAgent [Wang2024VideoAgentLV]. Our method shows more-stable performance.
5 Conclusion
We proposed an approach for Long-form Video Question Answering (LVQA) that achieves state-of-the-art results across 3 benchmarks datasets. Our propsoed Hierarchical Keyframe Selector and Sequential Visual LLM modules demonstrate the usefulness of keyframe selection and sequence-aware frame captioning. We also highlight the zero-shot capability for long-form video understanding of our LVNet framework requiring zero video level training, and our experiments demonstrate its clear advantage over prior work.