### v4

# Too Many Frames, Not All Useful: Efficient Strategies for Long-Form Video QA

Abstract
Long-form videos that span across wide temporal intervals are highly information redundant and contain multiple distinct events or entities that are often loosely related. Therefore, when performing long-form video question answering (LVQA), all information necessary to generate a correct response can often be contained within a small subset of frames. Recent literature explore the use of large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language. Such VLMs often independently caption a large number of frames uniformly sampled from long videos, which is not efficient and can mostly be redundant. Questioning these decision choices, we explore optimal strategies for key-frame selection that can significantly reduce these redundancies, namely Hierarchical Keyframe Selector. Our proposed framework, LVNet, achieves state-of-the-art performance at a comparable caption scale across three benchmark LVQA datasets: EgoSchema, NExT-QA, IntentQA. The code can be found at https://github.com/jongwoopark7978/LVNet
1 Introduction
Video understanding is a long-standing vision problem [1] with numerous real-world applications. It has been traditionally studied even before the era of differentiable representation learning, with hierarchical approaches focusing on longer videos [3, 15, 33, 13, 32]. Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering that involve generating human-style conversations in response to questions regarding videos [34, 51, 44].
The recent popularity of vision-language models (VLMs), particularly approaches connecting large language models (LLMs) to vision architectures [23, 21, 7], has resulted in significant improvements across visual question answering (VQA) tasks. These models demonstrate exceptional performance within the image domain, and their video variants [47, 28, 27] perform similarly on shorter videos, yet demonstrating limited performance on long-form video benchmarks [24, 16, 30, 31]. This can be attributed to the nature of long-form video benchmarks, which require both temporal sequence awareness and causal reasoning. An alternate line of works [52, 36, 16, 37] adapt LLMs that contain strong reasoning abilities for this task, using image VLMs to generate per-frame natural language descriptions, followed by video question answering purely within the language domain. However, these methods employ expensive VLMs to caption a large number of uniformly sampled frames. Such a design choice leading to high compute expense, is questioned in [4, 30, 37], and is the key motivation for our exploration of key frame selection, i.e. identifying a minimal set of frames most useful for correctly answering a given video-question pair.
Figure 1: (left) Our proposed LVNet contains a novel Hierarchical Keyframe Selector module to select keyframes and followed by VLM and LLM for caption and answer generation. Hierarchical Keyframe Selector initially begins by processing dense frames and keywords with the lighter module and progressively exploits heavier and more performance-oriented modules on a small set of frames, resulting from the reduction of keyframe candidates, to ensure efficient computation. (right) LVNet achieves state-of-the-art performance on the EgoSchema subset while utilizing only a fraction of captioned frames. In particular, LVNet obtains its highest accuracy of 68.2% with 12 captions, outperforming VideoAgent, a model using a similar-scale captions, by 8%. A more detailed analysis of accuracy vs. the number of captions is provided in 4.3.
Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) that performs efficient key-frame selection followed by VLM and LLM for caption and answer generation as illustrated in Figure 1. Aligned with prior work [52, 39, 37], the per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair. The scope of this work focuses on optimizing the prior two stages.
We summarize our key contributions as follows:
  • Novel Hierarchical Keyframe Selector module which are composed of three submodules: Temporal Scene Clustering (TSC), Coarse Keyframe Detector (CKD), and Fine Keyframe Detector (FKD).
    • Temporal Scene Clustering performs non-uniform frame sampling by clustering visually similar frames, minimizing redundancy in long videos and capturing key scenes. The lightweight module is used to efficiently filter dense frames.
    • Coarse Keyframe Detector generates keywords, representing atomic activities, using the given query and LLM. It assigns confidence scores to frames based on keyword relevance, sub-sampling high-confidence frames for improved interpretability over visual-only selection.
    • Fine Keyframe Detector refines frame selection by combining multiple frames using visual templating and Visual Language Model (VLM), enabling higher-level reasoning and natural language output for improved accuracy over the CKD’s keyword-based selection.
  • Long-form Video Understanding framework requiring zero video-level training.
Our proposed LVNet achieves state-of-the-art results compared to the models that utilize a similar number of captions on three long-form video question answering benchmarks—EgoSchema[24], NExT-QA[41], and IntentQA[22] as described in 4.2. This demonstrates the strong performance and general applicability of our approach.
2 Related Work
Video Question Answering: Visual question answering (VQA), a relatively recent task, involves generating free-form and open-ended textual content conditioned on an image and natural language text [2]. Its video variant, Video-VQA [49] replaces images with videos. While multiple early datasets focus on querying objects or events based on referential and spatial relations [44, 51, 49], later tasks require explicit temporal understanding of sequential events [18, 19, 50]. More recent datasets focus on longer videos containing multiple actions and scenes spread over a wide time interval (termed long-form videos) [41, 20] with questions additionally requiring causal reasoning to correctly answer. Referred to as long-form video question answering (LVQA), these benchmarks are constructed to specifically test strong causal and temporal reasoning [41] over long temporal windows [24]. Some works tackling such casual video VQA tasks leverage graph networks to model cross object / event relations [14, 42, 43]. A more recent line of works integrate large-language models (LLMs) to tackle this task [52, 36, 16, 37, 30, 39, 9] utilizing the strong reasoning skills of the LLMs. A common aspect is the use of a vision langauge model (VLM) to convert the frame level visual information into natural language. This is in turn input to the LLM which makes a final prediction.
Unlike these methods, our LVNet incorporates a unique Hierarchical Keyframe Selector that progressively reduces the number of keyframe candidates. Lighter modules are applied to dense frames, while heavier, more performance-focused modules are applied to a small set of frames for efficient computation. Additionally, LVNet does not require video-level training, unlike earlier supervised approaches.
Frame Selection in Videos: The task of frame selection in videos has been long explored in video [8, 53] with more recent works focused directly on long-form video question answering [4, 30]. Most similar to our work is [37] which employs an LLM based strategy for video frame selection. However, our proposed Hierarchical Keyframe Selector module differs with unique visual only operations as well as cross-modal operations handling much longer frame sequences extending up to 900 frames.
3 Method
In this section, we present our training-free (i.e. zero-shot) framework for long-form video QA, LVNet. Videos are a dense form of data with even a few seconds long clip being composed of 100s of frames (individual images). In the case of long-form videos, this frame count is even greater. However, the information necessary to answer a given question is often contained in a handful of those frames. Our framework tackles this challenge of selecting an optimal and minimal set of informative frames. We refer to this as keyframe selection. Given such a set of useful frames, we also establish optimal strategies for extracting their information using modern large language models (LLMs), taking into account their sequential nature.
Our proposed LVNet comprises of three components: a Hierarchical Keyframe Selector (HKS), a Vision Language Model (VLM), and a Large Language Model (LLM) as illustrated in Figure 1. The HKS, an efficient, hierarchical keyframe selector, is the core contribution of our work. First, the model processes 900 uniformly sampled frames and clusters them into distinct scenes using ResNet-18, a lightweight module that measures visual correspondence between frames. Next, it extracts keywords from a given natural language query via LLM and selects the frames most relevant to those keywords using CLIP-B/16. Finally, the selected frames are described in natural language by a more powerful and computationally intensive VLM. Finally, an LLM processes the language descriptions of the selected frames to answer a given query.
3.1 Background
Recent approaches utilizing LLMs for long video question answering (LVQA) [52, 36, 16, 30, 37] can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation. Note that the complexity of each stage varies across methods given their focus on different aspects of the LVQA task (e.g. frame selection in some is simply uniform sampling). In our work, we also follow this structure, but we focus on improving the frame selection stage. Under such a framework, our proposed HKS can serve as plug-in modules to replace the frame selection stage and the later two stages are similar to these prior works.
Figure 2: Detailed architecture of Hierarchical Keyframe Selector: Input frames first go through a Temporal Scene Clustering (TSC) mechanism, which divides the long-video into scenes, enabling per-scene subsampling. Next, a Coarse Keyframe Detector (CKD) is applied to select frames best-aligned with keywords relevant to the query. Finally, a Fine Keyframe detector further subsamples frames, by refining keyword alignements through a templated visual prompting.
3.2 Architecture
Consider a long video, $\mathbf{x}\in\mathbb{R}^{T\times C\times H\times W}$ with T, C, H, W for frames, channels, height, and width respectively. Its paired natural language query is referred as $\mathbf{q}$. Also consider a frame in $\mathbf{x}$ at timestamp $t$ as $\mathbf{x}[t]\in\mathbb{R}^{C\times H\times W}$. Our goal is to output the suitable response to this query based on information contained in the video. We refer this correct response as $\mathbf{r}$.
Our LVNet processes a given video-query ($\mathbf{x},\mathbf{q}$) pair to output a response, $\mathbf{\hat{r}}$. The HKS module initially processes this video-query pair, selects $T^{\prime}$ keyframes, and outputs a deterministically sub-sampled video $\mathbf{x^{\prime}}\in\mathbb{R}^{T^{\prime}\times C\times H\times W}$. Each of these $T^{\prime}$ frames is then passed through the captioning stage of our VLM to generate a set of natural language descriptions, $D=\{d_{1},d_{2},...d_{T^{\prime}}\}$ where $d_{i}$ describes the frame $\mathbf{x^{\prime}}[i]$. Finally, the LLM processes all descriptions $D$ and the query $\mathbf{q}$ to generate response $\mathbf{\hat{r}}$.
Our unique contributions here are the Hierarchical Keyframe Selector (HKS) and optimal prompts and description templates for the LLM when generating the final response which is also important as suggested by our experiments. The overall architecture is given in Figure 1 (left).
3.3 Hierarchical Keyframe Selector
Figure 3: Qualitative example: We illustrate a challenging long-video QA scenario from EgoSchema [24]. We consider an input of 900 frames, which first get clustered into scenes and subsampled to retain around 390 frames. Next, the Coarse Keyframe Detector selects only 32 frames out of them, based on the alignment with keywords (Here, keywords are extracted based on answer options, via an LLM). Such coarse keyframes are then ranked based on the combination of confidence value and temporal span, and grouped into four sets, each containing eight frames. These sets are then processed through visual templating (i.e. simple concatenation across space) and fed into a VLM for Fine Keyframe Detection, resulting in just 12 frames.
We now describe our proposed Hierarchical Keyframe Selector (HKS) module. As illustrated in Figure 2, our HKS comprises of three submodules, namely 1) Temporal Scene Clustering (TSC), 2) Coarse Keyframe Detector (CKD), and 3) Fine Keyframe Detector (FKD). The inputs to HKS contain $T$ frames with each of the three stages TSC, CKD, and FKD reducing this frame count to $T_{a}$, $T_{b}$, and $T_{c}=T^{\prime}$ respectively.
Temporal Scene Clustering (TSC): The role of TSC is to perform visual content aware preliminary frame sampling. The established approach for preliminary frame selection is uniform sampling. In contrast, TSC extracts per-frame visual features using a deep neural network followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets. Within each of the n sets, we uniformly sample $\leq\tau$ frames obtaining a total of $T_{a}\leq\tau\times n$. Our iterative clustering procedure is outlined in Algorithm 1. It calculates pairwise distances between all frames accounting for intra-frame local information using the extracted per-frame features, followed by n iterative frame similarity based grouping operations.
Our reasoning for TSC is that videos, especially long-form ones, contain extensive information redundancy across its temporal dimension. For example, in a long sub-sequence of mostly static frames from a single scene, beyond its first frame the rest may contain little additional information. We hypothesize that our TSC module will utilize frame level feature similarity to cluster the frames, such that each cluster corresponds to a common scene or motion. A single cluster could contain just one frame or significantly more based on frame feature similarities. This leads to a non-uniform sampling of frames across the entire video. In contrast to uniform sampling across the entire videos where each timestamp is treated equally, in our formulation the uniform sampling within segments will allow more frames to be sampled from information heavy temporal regions.
Coarse Keyframe Detector (CKD): Unlike the TSC in prior stage, CKD reasons across both visual and language modalities (using the paired textual query) to further sub-sample the $T_{a}$ frames output from the TSC module. The inputs to this CKD stage are the $T_{a}$ frames and the paired query $\mathbf{q}$. The module generates a set of keywords $K_{\text{CKD}}$, selects $T_{b}$ frames conditioned on these keywords, and outputs the selected $T_{b}$ frames with a confidence value and keyword assigned to each frame.
The CKD contains three elements: a keyword generation strategy, a dual-encoder image-text model, and an iterative algorithm for similarity based confidence assignment. The keyword generation strategy utilizes the given query, q alongside a hand-crafted templating operation or an LLM to select or generate suitable keywords. The dual-encoder image-text model uses a spatially aware contrastive langauge image pre-training (CLIP) network from [29]. For confidence assignment, we construct an algorithm as outlined in Algorithm 2 which processes two lists, one of frames and one of keywords, and then calculates their pairwise likelihood of occurrence to assign each frame a confidence value (that reflects its usefulness to answer the query, $\mathbf{q}$).
The keyword generation strategy converts the query $\mathbf{q}$ into a set of keywords, $K_{\text{CKD}}$ where the number of keywords $|K_{\text{CKD}}|\in[8,25]$. The dual-encoder model converts the keyword set $K_{\text{CKD}}$ and the frame set (containing $T_{a}$ frames) into two sets of embeddings. The embeddings of each set are comparable to each other using a cosine similarity distance metric given the training formulation of the CLIP model. These two embedding sets are processed by our confidence assignment algorithm in Appendix A.2 which assigns each of the $T_{a}$ frames a confidence score and a keyword. Note that the same keyword can be assigned to multiple frames. However, given the much higher frame count compared to the keyword count and the cross-frame redundancy, we do not assign multiple keywords to a single frame.
The key role of this module is language conditioned frame selection. For a single query, there can be multiple regions in a video that are highly informative but not useful or relevant in answering that query. A single query can also contain multiple different concepts and attributes that must be given attention to construct a correct answer: the keyword generation attempts to capture each of these distinct attributes. On the visual modality, a single frame will also encode multiple concepts and attributes. Our design choice for the spatially aware CLIPpy dual-encoder VLM from [29] is motivated by this nature of individual frames. Finally, confidence assignment takes into account these multiple modes of information within each frame and the query and suitably assigns a confidence score to each frame that reflects its relevance to the query. We also highlight how the confidence scores are directly linked to the related keyword (i.e. reason that makes the frame relevant), leading to better interpretability and the ability to perform further keyword-based refinement in later stages.
Fine Keyframe Detector (FKD): In the prior CKD stage, cross-modal selection utilizes a dual-encoder VLM that is constrained by the set of keywords provided and performs limited reasoning at frame level. In contrast, FKD uses a visual templating module to combine multiple frames and uses VLM to generate open-ended natural language output through higher-level reasoning. The input in this stage is the set of $F_{b}$ frames, with each frame having an assigned confidence score and keyword.
Our visual templating module partitions the $T_{b}$ frames into sets of 8 ordered by their confidence scores, arranges frame sets as grids to form a collage-style image, and annotates that image with visually identifiable tags corresponding to each frame. We further illustrate this process in Figure 3 (see Visual Templating column). Each of these visual templated images also contain a subset of the keywords that correspond to their 8 images. These resulting visual templated images along with a prompt containing their associated keywords and instructions to select a frame subset based on valid association between keywords and images (see Appendix A.3 for details) are input to the VLM. The output of the VLM is used to select a subset of each 8 image group. These frames are collected as the output of the FKD stage, overall resulting in $T_{c}$ frames.
The purpose of the initial visual templating module is to allow reasoning across a set of frames using the image-text VLM (which is trained to process a single image at time). This partitioning of the input $T_{b}$ frames is performed based on confidence scores from the prior stage and timestamps. The eight frames with top confidence scores are grouped into the first visual template, followed by the next eight and so forth. This ensures the VLM selects both high confidence concepts and low confidence concepts, accounting for biases and weaknesses in our CKD stage. After that, we temporally reorder some image sets with low confidence scores to cover keyframes distributed across long-range segments, while the sets with high confidence scores concentrate on keyframes in short-range segments. A total of 16 low-score frames are temporally reordered in this process. The algorithm is described in Algorithm 3 and the prompting technique is explained in Appendix A.3. Our intuition is that such a mechanism allows one to best utilize the complementary strengths of two different VLMs from CKD and FKD stages for better frame selection overall.
3.4 Comparison with Other Keyframe Selection Methods:
We aim to highlight the main advantage of the Hierarchical Keyframe Selector over other existing keyframe selection methods. Models like VideoAgent and TraveLER provide useful comparisons, as they work with a similar-scale number of captions. VideoAgent and TraveLER rely on uniform frame selection in the first iteration without analyzing the entire video. They identify important segments based solely on these initial frames and the LLM’s response, which can be problematic if the initial uniformly selected frames are not representative of the entire video or if the LLM misinterprets the captions and prompts. In such cases, the LLM might incorrectly identify segments for further analysis. If the LLM fails to pinpoint the correct segment initially, the entire process can break down because subsequent frames will be similar to the first set, leading the LLM to continuously select frames within or near the initial segment. Additionally, for videos that are as challenging or more difficult than EgoSchema in terms of temporal complexity and activities, these models may require numerous iterations to finalize keyframes selection. This results in higher computational and latency costs, as it necessitates numerous runs of resource-intensive VLM and LLM models.
In contrast, our method analyzes the entire video with high frame rates using a lightweight ResNet-18 [12] and segments the video non-uniformly based on scene continuity. We then select several frames in each segment by measuring feature similarity between frame features and keywords using the CLIP-B/16 (0.12B) [29] which is lighted than VideoAgent’s EVA-CLIP-8Bplus (8B). By reviewing the entire video and non-uniformly selecting keyframes based on scene continuity and similarity scores, these keyframes accurately represent the question-based important frames distribution in the entire video. Furthermore, we use VLM for a fine-grained selection of keyframes, improving keyframe selection when CLIP-B/16 struggles to understand detailed atomic activities in the frames. By hierarchically segmenting the video with different modules, the resulting segments and keyframes are more reliable than those from VideoAgent. Even with more challenging videos, our process only needs to go through the video once to collect keyframes, maintaining computational efficiency.
4 Experiments
In this section, we first discuss our experimental setup followed by quantitative evaluations comparing to existing baselines and ablations of our proposed components. We then present qualitative results for our method and outline some limitations of our approach.
4.1 Experimental Setup
Datasets: Given the training free nature of our framework, we do not utilize any video datasets for training. Datasets are used purely for evaluation. We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema [24], NExT-QA [41], and IntentQA [22]. The first dataset, EgoSchema, is a very long-form video question-answering dataset with 5031 questions. For each question, EgoSchema requires the correct answer to be selected between five given options based on a three-minute-long video clip. The second dataset, NExT-QA, is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer. These questions are further classified as Causal (Cau.), Temporal (Tem.), and Descriptive (Des.) and we evaluate on its validation set containing 4996 questions over 570 videos. The third dataset, IntentQA [22] is based on NExT-QA videos corresponding to temporal and causal reasoning quetions. It consists of 16k multiple-choice questions which are classified as Why?, How? or Before/After (B./A.).
Model Choices: For the HKS module, we use the ResNet-18 [12] for the TSC, CLIP-B/16 [29] for the CKD and GPT-4o for the FKD. We select ResNet-18 and CLIP-B/16 due to their smaller models sizes—0.01B and 0.12B, respectively–which are significantly lighter compared to GPT-4o, whose model size is expected to be on the scale of 100B-1T. This makes them well-suited for filtering dense frames efficiently. In line with previous state-of-the-art work [39, 52, 36], we employ GPT API, especially GPT-4o, for both VLM and LLM. This choice is driven by its cost-effectiveness and lighter computational requirements compared to GPT-4. GPT-4o is used as the VLM for generating captions and as the LLM for answering questions in our framework.
Table 1: Combined Results on EgoSchema [24], NExT-QA [41], and IntentQA [22]. We compare LVNet against prior zero-shot models across three datasets, highlighting different task splits. The models are ordered based on the number of captions processed before answering the question. LVNet achieves state-of-the-art accuracies of 71.7%, 61.1%, and 72.9% on the three datasets, respectively, using just 12 frames compared to the models using the similar number of captions. Models with video-caption pretraining or utilizing significantly more captions than the 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fairness with image-level pretraining or highlight caption efficiency.
Model EgoSchema NExT-QA IntentQA
Cap. Acc. (%) Cap. Cau. (%) Tem. (%) Des. (%) All (%) Cap. Why? (%) How? (%) B./A. (%) All (%)
Vamos [36] - 48.3 - - - - - - - - - -
IG-VLM [17] - 59.8 - 69.8 63.6 74.7 68.6 - - - - 65.3
VideoLLaMA 2 [5] - 53.3 - - - - - - - - - -
InternVideo2 [38] - 60.2 - - - - - - - - - -
Tarsier [35] - 61.7 - - - - 79.2 - - - - -
VIOLET [10] 5 19.9 - - - - - - - - - -
mPLUG-Owl [46] 5 31.1 - - - - - - - - - -
VideoAgent [37] 8.4 54.1 8.2 72.7 64.5 81.1 71.3 - - - - -
MVU [30] 16 37.6 16 55.4 48.1 64.1 55.2 - - - - -
MoReVQA [25] 30 51.7 16 70.2 64.6 - 69.2 - - - - -
TraveLER [37] - - $\sim$25 70.0 60.5 78.2 68.2 - - - - -
VFC [26] - - 32 45.4 51.6 64.1 51.5 - - - - -
SeViLA [48] 32 22.7 32 61.3 61.5 75.6 63.6 32 - - - 60.9
ProViQ [6] 60 57.1 60 - - - 64.6 - - - - -
VideoTree [40] 63.2 61.1 $\leq$56 75.2 67.0 81.3 73.5 $\leq$56 - - - 66.9
FrozenBiLM [45] 90 26.9 - - - - - - - - - -
LifelongMemory [39] 90 62.1 - - - - - - - - - -
LangRepo [16] 180 41.2 90 64.4 51.4 69.1 60.9 90 62.8 62.4 47.8 59.1
LLoVi [52] 180 50.3 90 69.5 61.0 75.6 67.7 90 68.4 67.4 51.1 64.0
LVNet (ours) 12 61.1 12 75.0 65.5 81.5 72.9 12 75.0 74.4 62.1 71.7
4.2 Evaluation
Quantitative Results:
We evaluate LVNet on the EgoSchema, NExT-QA, and IntentQA dataset and present our results in Table 1. Models with video-caption pretraining are de-emphasized in grey to ensure fairness with image-level pertaining. Models utilizing significantly more captions than the 12 frames are downplayed in light green to consider caption efficiency. For EgoSchema, we achieve 61.1% on the fullest, the highest among the models utilizing approximately 12 captions. This result outperforms VideoAgent, the next best model using 8.4 captions, by 7%. To ensure a fair comparison, we ablate the number of captions for LVNet in LABEL:tab:ablation:numberofframecaptions, observing that LVNet with 8 captions still outperforms VideoAgent with 8.4 captions by 4.2% in the EgoSchema subset. The superior accuracy of LVNet over other keyframe selectors, such as those used in VideoAgent and TraveLER, is discussed in Section 3.4.
We next evaluate on the NExT-QA[41] dataset. This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs. Our approach achives state-of-the-art performance on this benchmark outperforming prior work among the models utilizing approximately 12 captions. In fact, our LVNet with just 12 frame captions achieves 72.9% overall accuracy, outperforming VideoAgent [37] (71.3% at 8.2 captions) by 1.6%.
We finally evaluate on the IntentQA[22] dataset. Our approach achieves state-of-the-art results on this benchmark, outperforming all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames. In fact, LVNet shows a substantial improvement (+4.8%) over the next best model, VideoTree [40] (66.9% with 90 captions), achieving 71.7% accuracy with only 12 frame captions.
Given the generative nature of VQA tasks as well as the limited availability and noisy nature of fully-annotated video VQA corpora, building generalizable fully-supervised models are challenging for these tasks. Nevertheless, we highlight how our zero-shot and video level training-free framework is competitive with the best supervised approaches on this dataset. This indicates the promise of utilizing pretrained models, especially those equipped with extensive world knowledge and reasoning skills from alternate modality specific learning (i.e. in our cases image domain VLMs and language domain LLMs).
Qualitative Analysis of the Hierarhical Keyframe Selector:
We compare the open-ended responses of LVNet and the uniform sampling method in Figure 4 to understand the effectiveness of the hierarchical keyframe selector in LVNet. The frames chosen by LVNet and the naive uniform sampling method are indicated by blue and red checkmarks in the images, respectively. LVNet selects frames at 5, 69, and 135 seconds by executing the hierarchical keyframe selector and generates captions based on those frames. When we feed the concatenated captions to the LLM to answer the given question: "Based on the video, what are the three main types of tools that C uses…" in an open-ended manner, the output identifies two main activities: welding torches and measuring tapes, among the three main activities described in Option 3 (welding handle, hammer, and measuring tape), which is the correct answer. This leads LVNet to choose the correct option.
In contrast, the uniform sampling method selects frames at 0, 16, and 32 seconds and generates captions based on those frames. Similarly, when we feed the concatenated captions to the LLM to answer the same question, the output identifies only one activity—welding tools—resulting in the selection of the incorrect option. This example highlights the importance of keyframe selection and demonstrates the effectiveness of hierarchical keyframe selection in LVNet.
Figure 4: Open-ended Responses from LVNet vs Uniform Sampling: The frames chosen by LVNet and the naive uniform sampling method are indicated with blue and red checkmarks, respectively. LVNet identifies both welding torches and measuring tapes, choosing the correct option, whereas uniform sampling only detects welding tools and selects the incorrect answer. The blue, red, and purple highlights correspond to the three main activities in the video—welding a handle, using a hammer, and using a measuring tape, respectively.
4.3 Ablations
Table 2: Ablations on EgoSchema [24]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot video VQA.
Visual Templating order Acc. (%)
Temporal 65.2
Hybrid (Confidence + Temporal) 68.2
TSC CKD FKD Acc.
68.2
65.8
64.5
62.6
Frames 6.4 8 8.4 11 12 16
VideoAgent[37] 58.4 - 60.2 57.4 - -
LVNet (ours) - 64.4 - - 68.2 67.8
In this section, we present ablations on key design decisions such as the sorting order in FKD, the number of frames for captions, and the effect of different components in HKS. In all ablations, we use a subset of EgoSchema [24], composed of 500 videos. Additional ablations about Choice of LLM and Effect of Patch Size on Keyword Matching in CKD are in Appendix A.1
Visual Templating Order:
In visual templating, prioritizing frames by keyword confidence scores followed by reordering low-confidence frames based on timestamp is more effective than relying solely on temporal order. In this hybrid approach, high-confidence frames are selected from shorter segments by sampling three keyframes per set of eight, grouped by confidence rather than timespan. Conversely, low-confidence keyframes, which are crucial but more visually challenging for keyword matching, are sampled from broader video sections, with three frames per set of eight grouped by timestamp. This hybrid approach outperforms purely temporal ordering by 3%.
Number of Frame Captions:
We conducted an ablation study on the number of frame captions used in our setup, comparing it to VideoAgent [37], which operates with a similarly small number of captions. Our findings show that LVNet achives its highest accuracy 68.2% with 12 captions, while even its lowest accuracy 64.4% with 8 captions still surpasses VideoAgent’s best accuracy 60.2% with 8.4 captions.
Effect of Hierarchical Keyframe Modules:
This table demonstrates the impact of incrementally adding the temporal scene clustering (TSC), coarse keyframe detector (CKD), and fine keyframe detector (FKD) modules. Without any of these modules, the model relies on uniform sampling and achieves 62.6%. When TSC is added and 12 frames are selected uniformly, the accuracy increases to 64.5%. Adding both TSC and CKD raises the accuracy to 65.8%. Finally, incorporating all three modules—TSC, CKD, and FKD—into the model, which is LVNet, results in an accuracy of 68.2%. This demonstrates the importance of including all modules in LVNet for optimal performance.
5 Conclusion
We proposed a novel approach for Long-form Video Question Answering (LVQA) that achieves state-of-the-art performance compared to the model using the similar-scale captions across 3 benchmarks datasets. Our Hierarchical Keyframe Selector demonstrates the effectiveness of keyframe selection in understanding a very long-form video QA. Additionally, we highlight the zero-shot capability for long-form video comprehension of our LVNet framework, which requires no video-level training. Our experiments showcase its significant advantage over previous methods.
Acknowledgements:
This work was supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government [24ZB1200, Research of Human-centered autonomous intelligence system original technology]. This work was also supported by the Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No. RS-2024-00336738, Development of Complex Task Planning Technologies for Autonomous Agents, 100%)
Appendix
Appendix A.1 Additional Ablations
In this section, we present additional experiments conducted to inform the LVNet’s design. We have tested different LLMs and experimented with various scales of the visual feature map.
Table A.3: Ablations on EgoSchema [24]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA.
LLM Acc. (%)
GPT-3.5 61.0
GPT-4 65.4
GPT-4o 68.2
Patch Size Acc. (%)
1x1 63.6
7x7 66.2
14x14 68.2
Choice of LLM:
Our results show that GPT-4o outperforms both GPT-4 and GPT-3.5 by 2.8% and 7.2%, respectively. Given that GPT-4o is more cost-effective and lightweight compared to GPT-4, we have selected it as our default LLM.
Effect of Patch Size on Keyword Matching in CKD:
Since keywords can represent activities spanning the entire image or confined to a small region, we adjust the resolution of the visual feature map output from the spatially aware contrastive image pre-training (CLIP) network [29] to match keywords. Our findings show that higher resolutions lead to better accuracy. In LVNet, we use a 14$\times$14 feature map and determine the confidence level of the keyword by selecting the maximum value between the 14$\times$14 patches and the keyword’s text embedding.
Appendix A.2 Algorithms in Detail
Our algorithms are presented in full detail in Algorithm 1, Algorithm 2, and Algorithm 3. TSC in Algorithm 1 extracts per-frame visual features using ResNet-18, followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets. Within each of the $n$ sets, we uniformly sample $\leq\tau$ frames, obtaining a total of $T_{a}\leq\tau\times n$ frames. For example, LVNet sets $\psi=5,\lambda=12,\tau=18$, resulting in approximately $n\sim 25$ and $T_{a}\sim 390$ on the EgoSchema dataset. CKD in Algorithm 2 selects top $L$ frames based on similarity/confidence scores, which are calculated using cosine similarity between frames and keywords with CLIP-B/16. LVNet employs $L=32,len(K)\leq 25$ on the EgoSchema dataset. FKD in Algorithm 3 sorts frames and their corresponding keywords by confidence scores, and reorder the $K$ frames with the lowest scores temporally. It groups frames sequentially into visual templates, each consisting of $N$ frames. From each template, the $M$ frames and keywords most relevant among the $N$ pairs are selected using GPT-4o. We set $L=32,K=16,N=8,M=3$.
1:Require: ResNet-18 [11] pretrained on imagenet dataset $f$, frame list $\textbf{List}_{frame}$, image index list $\textbf{List}_{index}\in\{1,\dots,N\}$, minimum number of list length $\psi$, temperature $\lambda$, number of sample $\tau$, function to find index of x in list w index(x, w), and function to sort list sort(List)
2:for all $img^{i}$ in $\textbf{List}_{frame}$ do
3:   $\mathbf{F}^{i}\leftarrow f(img^{i})$
4:   $\textbf{List}_{feat}$.insert($\mathbf{F}^{i}$)
5:end for
6:for all $\mathbf{F}^{i}$ in $\textbf{List}_{feat}$ do
7:   $\textbf{List}_{dist}\leftarrow\frac{\sum_{y}\sum_{x}{\sqrt{(\mathbf{F_{i}}-\textbf{List}_{feat})^{2}}}}{x\times y}$
8:   $\textbf{M}_{dist}$.insert($\textbf{List}_{dist}$)
9:end for
10:while length of $\textbf{List}_{index}>\psi$ do
11:   $\textbf{List}_{sample}\leftarrow\emptyset$
12:   $\textbf{List}_{\delta}\leftarrow\emptyset$
13:   $i\leftarrow\textbf{List}_{index}$.pop(00)
14:   $\textbf{p}^{i}\leftarrow\texttt{softmax}(\textbf{M}^{i}_{dist})$
15:   $\mu_{\mathbf{\textbf{p}^{i}}},\sigma_{\mathbf{\textbf{p}^{i}}}\leftarrow\texttt{mean}(\textbf{p}^{i}),\texttt{std}(\textbf{p}^{i})$
16:   $\beta\leftarrow\mu_{\mathbf{\textbf{p}^{i}}}-\sigma_{\mathbf{\textbf{p}^{i}}}\sum_{i=0}{e^{1-i/\lambda}}$
17:   for all prob in $\textbf{p}^{i}$ do
18:      if prob < $\beta$ then
19:         $\textbf{List}_{selected}$.insert($\textbf{index}(\texttt{prob},\textbf{p}^{i})$)
20:      end if
21:   end for
22:   for all $\gamma$ in $\textbf{List}_{selected}$ do
23:      $\delta$leftarrow $\gamma~{}th$ value in $\textbf{List}_{index}$
24:      $\textbf{List}_{\delta}$.insert($\delta$)
25:      $\textbf{List}_{index}$.pop($\gamma$)
26:   end for
27:   $\textbf{List}_{\delta}$.insert($i$)
28:   $\textbf{List}_{sample}$leftarrow sample $\tau$ items from $\textbf{List}_{\delta}$
29:   $\textbf{sort}(\textbf{List}_{sample})$
30:   for all $frame^{j}$ in $\textbf{List}_{frame}$ do
31:      if $j$ in $\textbf{List}_{sample}$ then
32:         Outputs.insert($frame^{j}$)
33:      end if
34:   end for
35:end while
Algorithm 1 Temporal Scene Clustering
1:Require: keyword set $\mathbf{K}$, image set $\mathbf{I}$, total length of selected image set $L$, function to calculate similarity matrix $\textbf{sim}(\mathbf{K},\mathbf{I})$, function to sort similarity matrix and return indices $\textbf{sort}(\mathbf{S})$
2:$\mathbf{S}\leftarrow\textbf{sim}(\mathbf{K},\mathbf{I})$
3:$\mathbf{S}_{sorted},\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
4:Initialize $\mathbf{P}_{best}$ as an empty list
5:Initialize $\mathbf{I}_{selected}$ as an empty set
6:while length of $\mathbf{I}_{selected}<L$ do
7:   for $k\in\mathbf{K}$ do
8:      for $i\in\mathbf{I}$ do
9:         $i_{index}\leftarrow\mathbf{idx}_{sorted}[k][i]$
10:         if $i_{index}$ not in $\mathbf{I}_{selected}$ then
11:            $\mathbf{P}_{best}$.insert($k,i_{index}$)
12:            $\mathbf{I}_{selected}$.insert($i_{index}$)
13:            break
14:         end if
15:      end for
16:      if length of $\mathbf{I}_{selected}\geq L$ then
17:         break
18:      end if
19:   end for
20:end while
21:return $\mathbf{P}_{best}$
Algorithm 2 Keyword-Image Matching Process in CKD
[…]
2:$\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
[…]
4:$\mathbf{idx}_{temporal}\leftarrow\textbf{temporal\_order}(\mathbf{idx}_{low\_sim})$
[…]
6:$\mathbf{I}_{ordered},\mathbf{K}_{ordered}\leftarrow\mathbf{I}[\mathbf{idx}_{final}],\mathbf{K}[\mathbf{idx}_{final}]$
[…]
8:for each $\mathbf{set}\in\mathbf{sets}$ do
9:   $\mathbf{I}_{selected}\leftarrow\texttt{select\_top\_M}(\mathbf{set},M)$
10:end for
11:return $\mathbf{I}_{selected}$
Algorithm 3 Fine Keyframe Detection Process (FKD)
Appendix A.3 Prompting: Fine Keyframe Detector
We prompt the VLM to select frames that are most compatible with the list of given keywords. Each template image contains 8 images, and their order is described in language (e.g. top left to right, bottom left to right) and the VLM outputs the selected images according to our prompting as described in Figure A.5.
Figure A.5: Prompt for Fine Keyframe Detection: The figure illustrates the input image, the prompt provided to the VLM, and the output. The input image represents a visual template composed of eight frames, and the prompt requests the three best frames along with their corresponding keywords. The output displays the top three selected frames and their associated keywords.