### v5

# Too Many Frames, Not All Useful: Efficient Strategies for Long-Form Video QA

Abstract
Long-form videos that span across wide temporal intervals are highly information redundant and contain multiple distinct events or entities that are often loosely related. Therefore, when performing long-form video question answering (LVQA), all information necessary to generate a correct response can often be contained within a small subset of frames. Recent literature explore use of large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language. Such VLMs often independently caption a large number of frames uniformly sampled from long videos, which is not efficient and can mostly be redundant. Questioning these decision choices, we explore optimal strategies for key-frame selection that can significantly reduce these redundancies, namely Hierarchical Keyframe Selector. Our proposed framework, LVNet, achieves state-of-the-art performance at a comparable caption scale across three benchmark LVQA datasets: EgoSchema, NExT-QA, and IntentQA, while also demonstrating a strong performance on videos up to an hour long in VideoMME. The code can be found at https://github.com/jongwoopark7978/LVNet
Too Many Frames, Not All Useful:
Efficient Strategies for Long-Form Video QA
Jongwoo Park 1  Kanchana Ranasinghe 1  Kumara Kahatapitiya 1 Wonjeong Ryu2  Donghyun Kim 3  Michael S. Ryoo 1 1Stony Brook University   2KAIST AI   3Korea University jongwopark@cs.stonybrook.edu
1 Introduction
Video understanding is a long-standing vision problem (Aggarwal and Ryoo, 2011) with numerous real-world applications. It has been traditionally studied even before the era of differentiable representation learning, with hierarchical approaches focusing on longer videos (Allen and Ferguson, 1994; Ivanov and Bobick, 2000; Shi et al., 2004; Hongeng et al., 2004; Ryoo and Aggarwal, 2006). Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering that involve generating human-style conversations in response to questions regarding videos (Tapaswi et al., 2016; Zeng et al., 2017; Xu et al., 2017).
Figure 1: LVNet achieves state-of-the-art performance on EgoSchema subset while utilizing only a fraction of captioned frames. In particular, LVNet obtains its highest accuracy of 68.2% with 12 captions (VLM calls), outperforming VideoAgent and VideoTree, models using a similar-scale captions, by +8% and +5.7% (more details in Section 4.3).
Recent popularity of vision-language models (VLMs), particularly approaches connecting large language models (LLMs) to vision architectures (Liu et al., 2023; Li et al., 2023b; Dai et al., 2023), has resulted in significant improvements across visual question answering (VQA) tasks. These models demonstrate exceptional performance within the image domain, and their video variants (Yu et al., 2023; Papalampidi et al., 2023; Maaz et al., 2023) perform similarly on shorter videos, yet demonstrate limited performance on long-form video benchmarks (Mangalam et al., 2023; Kahatapitiya et al., 2024; Rawal et al., 2024). This can be attributed to the nature of long-form video benchmarks, which require both temporal sequence awareness and causal reasoning. An alternate line of works (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Wang et al., 2024b) adapt LLMs that contain strong reasoning abilities for this task, using image VLMs to generate per-frame natural language descriptions, followed by video question answering purely within the language domain. However, these methods employ expensive VLMs to caption a large number of uniformly sampled frames. Such a design choice leading to high compute expense, is questioned in (Buch et al., 2022; Ranasinghe et al., 2024; Wang et al., 2024b), and is the key motivation for our exploration of key frame selection, i.e. identifying a minimal set of frames most useful for correctly answering a given video-question pair.
Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) that performs efficient key-frame selection followed by VLM and LLM for caption and answer generation as illustrated in Fig. 2. Aligned with prior work (Zhang et al., 2023; Wang et al., 2024d, b), the per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair. As shown in Fig. 1, LVNet achieves strong performance using a small set of keyframes from the HKS. The scope of this work focuses on optimizing the prior two stages. We summarize our key contributions as follows:
1. Hierarchical Keyframe Selector (HKS): The proposed HKS consists of three submodules for efficient keyframe selection.
  1. (a)
    Temporal Scene Clustering (TSC)
    • Performs non-uniform frame sampling by clustering visually similar frames.
    • Reduces redundancy in long videos while capturing key scenes.
    • A lightweight module for efficient filtering of dense frames.
  2. (b)
    Coarse Keyframe Detector (CKD)
    • Generates keywords representing atomic activities using the given query and an LLM.
    • Assigns confidence scores to frames based on keyword relevance.
    • Samples high-confidence frames for improved interpretability over visual-only selection.
  3. (c)
    Fine Keyframe Detector (FKD)
    • Refines frame selection by combining multiple frames using visual templating and a VLM.
    • Enables higher-level reasoning and natural language-based selection.
    • Achieves better accuracy than CKD’s keyword-based selection.
2. Zero-Shot Long-Form Video Understanding: Our framework operates zero-shot without requiring video-level training. This makes it highly efficient for long-form video understanding.
Figure 2: (top) Overview: LVNet uses a Hierarchical Keyframe Selector (HKS) module to select keyframes, followed by VLM & LLM for caption and answer generation. (below) HKS Module processes dense frames with lighter modules and progressively exploits heavier, more performance-oriented modules on smaller subsets of frames to ensure efficient computation.
Proposed LVNet achieves state-of-the-art results compared to models utilizing similar number of captions on three long-form video question answering benchmarks— EgoSchema, NExT-QA, and IntentQA(Sec. 4.2). This demonstrates strong performance and generality of our approach.
2 Related Work
Video Question Answering: Visual question answering (VQA) involves generating open-ended textual content conditioned on an image and natural language query (Agrawal et al., 2015). Its video variant, Video-VQA (Yu et al., 2019a) replaces images with videos. Multiple early datasets focus on querying objects or events based on referential and spatial relations (Xu et al., 2017; Zeng et al., 2017; Yu et al., 2019a). Later tasks require explicit temporal understanding of sequential events (Lei et al., 2018, 2020; Yu et al., 2019b). More recent datasets focus on longer videos containing multiple actions and scenes spread over wide time intervals (termed long-form videos) (Xiao et al., 2021; Li et al., 2022). Referred to as long-form video question answering (LVQA), these benchmarks are constructed to specifically test strong causal and temporal reasoning (Xiao et al., 2021) over long temporal windows (Mangalam et al., 2023). Some works tackling such video VQA tasks leverage graph networks to model cross object / event relations (Hosseini et al., 2022; Xiao et al., 2022a, b). A more recent line of works integrate LLMs to tackle this task (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Wang et al., 2024b; Ranasinghe et al., 2024; Wang et al., 2024d; Fan et al., 2024) utilizing the strong reasoning skills of LLMs. A common aspect is the use of a vision language model (VLM) to convert frame level visual information into natural language. This in turn is input to the LLM which makes a final prediction.
Unlike these methods, LVNet incorporates a unique Hierarchical Keyframe Selector that progressively reduces the number of keyframe candidates. Lighter modules are applied to dense frames, while heavier, more performance-focused modules are applied to a small subset of filtered frames. Additionally, LVNet does not require video-level training, unlike earlier supervised approaches.
Feature Ours VA Tr. VT
(effective selection)
Uses non-uniform sampling
Scene continuity-based selection
Robust to initial frames
Fine-grained visual refinement
(compute efficient)
Lightweight feature extraction
Single pass inference
Table 1: LVNet exhibits unique features compared to VideoAgent (VA) Fan et al. (2024), Traveler (Tr.) Shang et al. (2024) and VideoTree (VT) Wang et al. (2024e). See Appendix A.5 for details.
Frame Selection in Videos: The task of frame selection in videos has been long explored in video (Davis and Bobick, 1997; Zhao et al., 2017) with more recent works focused directly on long-form video question answering (Buch et al., 2022; Wang et al., 2024e; Fan et al., 2024). Most similar to our work is (Wang et al., 2024b) which employs an LLM based strategy for video frame selection. However, our LVNet differs with several unique features as summarized in Table 1.
3 Method
In this section, we present our training-free (i.e. zero-shot) framework for long-form video QA, LVNet. Videos are a dense form of data with even a few seconds long clip being composed of 100s of frames (individual images). In the case of long-form videos, this frame count is even greater. However, the information necessary to answer a given question is often contained in a handful of those frames. Our framework tackles this challenge of selecting an optimal and minimal set of informative frames. We refer to this as keyframe selection. Given such a set of useful frames, we also establish optimal strategies for extracting their information using modern large language models (LLMs), taking into account their sequential nature.
Our proposed LVNet comprises of three components: a Hierarchical Keyframe Selector (HKS), a Vision Language Model (VLM), and a Large Language Model (LLM) as illustrated in Figure 2. The HKS, an efficient, hierarchical keyframe selector, is the core contribution of our work. First, the model processes 900 uniformly sampled frames and clusters them into distinct scenes Next, it extracts keywords from a given natural language query via LLM and selects the frames most relevant to those keywords. Finally, the selected frames are described in natural language by a more powerful and computationally intensive VLM. Finally, an LLM processes the language descriptions of the selected frames to answer a given query.
3.1 Background
Recent approaches utilizing LLMs for long video question answering (LVQA) (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Ranasinghe et al., 2024; Wang et al., 2024b) can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation. Note that the complexity of each stage varies across methods given their focus on different aspects of the LVQA task (e.g. frame selection in some is simply uniform sampling). In our work, we also follow this structure, but we focus on improving the frame selection stage. Under such a framework, our proposed HKS can serve as plug-in modules to replace the frame selection stage and the later two stages are similar to these prior works.
Figure 3: Qualitative example: We illustrate a challenging long-video QA scenario from EgoSchema (Mangalam et al., 2023). We consider an input of 900 frames, which first get clustered into scenes and subsampled to retain around 390 frames. Next, the Coarse Keyframe Detector selects only 32 frames out of them, based on the alignment with keywords (Here, keywords are extracted based on answer options, via an LLM). Such coarse keyframes are then ranked based on the combination of confidence value and temporal span, and grouped into four sets, each containing eight frames. These sets are then processed through visual templating (i.e. simple concatenation across space) and fed into a VLM for Fine Keyframe Detection, resulting in just 12 frames.
3.2 Architecture
Consider a video, $\mathbf{x}\in\mathbb{R}^{T\times C\times H\times W}$ with T, C, H, W for frames, channels, height, width respectively and its paired natural language query $\mathbf{q}$. Also consider a frame in $\mathbf{x}$ at timestamp $t$ as $\mathbf{x}[t]\in\mathbb{R}^{C\times H\times W}$. Our goal is to output a response, referred as $\mathbf{r}$, suitable for the given query $\mathbf{q}$ based on information contained in the video $\mathbf{x}$.
Our LVNet processes a given video-query ($\mathbf{x},\mathbf{q}$) pair to output a response, $\mathbf{\hat{r}}$. The HKS module initially processes this video-query pair, selects $T^{\prime}$ keyframes, and outputs a deterministically sub-sampled video $\mathbf{x^{\prime}}\in\mathbb{R}^{T^{\prime}\times C\times H\times W}$. Each of these $T^{\prime}$ frames is then passed through the captioning stage of our VLM to generate a set of natural language descriptions, $D=\{d_{1},d_{2},...d_{T^{\prime}}\}$ where $d_{i}$ describes the frame $\mathbf{x^{\prime}}[i]$. Finally, the LLM processes all descriptions $D$ and the query $\mathbf{q}$ to generate response $\mathbf{\hat{r}}$. We illustrate this overall architecture in Figure 2.
3.3 Hierarchical Keyframe Selector
We now describe our proposed Hierarchical Keyframe Selector (HKS) module. As illustrated in Figure 2, our proposed HKS comprises of three sequential submodules, each reducing the frame count to $T_{a}$, $T_{b}$, and $T_{c}=T^{\prime}$ respectively.
Temporal Scene Clustering (TSC): The role of TSC is to perform visual content aware preliminary frame sampling. The established approach for preliminary frame selection is uniform sampling (limited to at most 200 frames). In contrast, TSC processes 900 to 1800 uniformly sampled frames to extract per-frame visual features using a lightweight deep neural network (ResNet-18) followed by a clustering procedure to identify $n$ non-overlapping frame sets. Within each of the n sets, we uniformly sample $\leq\tau$ frames obtaining a total of $T_{a}\leq\tau\times n$. Our iterative clustering procedure is outlined in Algorithm 1. It calculates pairwise distances between all frames accounting for intra-frame local information using the extracted per-frame features, followed by n iterative frame similarity based clustering operations. A single cluster could contain just one frame or significantly more based on frame feature similarities, leading to a non-uniform sampling of frames across the entire video. This allows more frames to be sampled from the information heavy temporal regions of videos.
Coarse Keyframe Detector (CKD): Unlike TSC in the prior stage, CKD reasons across both visual and language modalities (using the paired textual query, $q$) to further sub-sample $T_{a}$ into $T_{b}$ frames. CKD contains three elements: keyword generation strategy, dual-encoder image-text model, and similarity based confidence assignment algorithm. Keyword generation utilizes the given query, q, alongside hand-crafted templating operations or an LLM to select or generate suitable keywords. The dual-encoder image-text model uses a spatially aware contrastive language image pre-training (CLIP) network from (Ranasinghe et al., 2023). For confidence assignment, we construct an algorithm as outlined in Algorithm 2 which processes two lists, one of frames and one of keywords, and then calculates their pairwise likelihood of occurrence to assign each frame a confidence value (that reflects its usefulness to answer the query, $\mathbf{q}$). See  Appendix A.3 for more details.
For a single query, there can be multiple regions in a video that are highly informative but not useful or relevant in answering that query. A single query can also contain multiple different concepts and attributes that must be given attention to construct a correct answer: the keyword generation attempts to capture each of these distinct attributes. On the visual modality, a single frame will also encode multiple concepts and attributes. Our design choice for the spatially aware CLIPpy dual-encoder VLM from (Ranasinghe et al., 2023) is motivated by this nature of individual frames. Finally, confidence assignment takes into account these multiple modes of information within each frame and the query to suitably assign confidence scores to each frame that reflects its query relevance. We also highlight how the confidence scores are directly linked to the related keyword (i.e. reason that makes the frame relevant), leading to better interpretability and the ability to perform further keyword-based refinement in later stages.
Fine Keyframe Detector (FKD): In the prior CKD stage, cross-modal selection utilizes a dual-encoder VLM that is constrained by the set of keywords provided and performs limited reasoning at frame level. In contrast, FKD uses a visual templating module to combine multiple frames and uses VLM to generate open-ended natural language output through higher-level reasoning. The input in this stage is the set of $F_{b}$ frames, with each frame having an assigned confidence score and keyword.
Our visual templating module partitions the $T_{b}$ frames into sets of 8 ordered by their confidence scores, arranges frame sets as grids to form a collage-style image, and annotates that image with visually identifiable tags corresponding to each frame. We further illustrate this process in Figure 3 (see Visual Templating column). Each of these visual templated images also contain a subset of keywords that correspond to their 8 images. These resulting visual templated images along with a prompt containing their associated keywords and instructions to select a frame subset based on valid association between keywords and images (see Appendix A.4 for details) are input to the VLM. The output of the VLM is used to select a subset of each 8 image group. These frames are collected as output of the FKD stage, overall resulting in $T_{c}$ frames.
The purpose of the initial visual templating module is to allow reasoning across a set of frames using the image-text VLM (which is trained to process a single image at time). This partitioning of the input $T_{b}$ frames is performed based on confidence scores from the prior stage and timestamps. The eight frames with top confidence scores are grouped into the first visual template, followed by the next eight and so forth. This ensures the VLM selects both high confidence concepts and low confidence concepts, accounting for biases and weaknesses in our CKD stage. After that, we temporally reorder some image sets with low confidence scores to cover keyframes distributed across long-range segments, while the sets with high confidence scores concentrate on keyframes in short-range segments. A total of 16 low-score frames are temporally reordered in this process. The algorithm is described in Algorithm 3 and the prompting technique is explained in Appendix A.4. Our intuition is that such a mechanism allows one to best utilize the complementary strengths of two different VLMs from CKD and FKD stages for better frame selection overall.
4 Experiments
In this section, we first discuss our experimental setup followed by quantitative evaluations comparing to existing baselines and ablations of our proposed components. We then present qualitative results for our method and outline some limitations of our approach.
Model EgoSchema NExT-QA IntentQA
Cap. Acc. (%) Cap. Acc. (%) Cap. Acc. (%)
Vamos (Wang et al., 2023) - 48.3 - - - -
IG-VLM (Kim et al., 2024) - 59.8 - 68.6 - 65.3
VideoLLaMA 2 (Cheng et al., 2024) - 53.3 - - - -
InternVideo2 (Wang et al., 2024c) - 60.2 - - - -
Tarsier (Wang et al., 2024a) - 61.7 - 79.2 - -
VIOLET (Fu et al., 2023) 5 19.9 - - - -
mPLUG-Owl (Ye et al., 2023) 5 31.1 - - - -
VideoAgent (Wang et al., 2024b) 8.4 54.1 8.2 71.3 - -
MVU (Ranasinghe et al., 2024) 16 37.6 16 55.2 - -
MoReVQA (Min et al., 2024) 30 51.7 16 69.2 - -
VFC (Momeni et al., 2023) - - 32 51.5 - -
SeViLA (Yu et al., 2024) 32 22.7 32 63.6 32 60.9
ProViQ (Choudhury et al., 2023) 60 57.1 60 64.6 - -
VideoTree (Wang et al., 2024e) 62.4 61.1 (56) 73.5 (56) 66.9
FrozenBiLM (Yang et al., 2022) 90 26.9 - - - -
LifelongMemory (Wang et al., 2024d) 90 62.1 - - - -
TraveLER (Shang et al., 2024) (101) 53.3 (65) 68.2 - -
LangRepo (Kahatapitiya et al., 2024) 180 41.2 90 60.9 90 59.1
LLoVi (Zhang et al., 2023) 180 50.3 90 67.7 90 64.0
LVNet (ours) 12 61.1 12 72.9 12 71.7
Table 2: Long Video Evalation: LVNet achieves state-of-the-art accuracies of 71.7%, 61.1%, and 72.9% on EgoSchema, NExT-QA, and IntentQA datasets respectively using just 12 frames compared to models using a similar number of captions. Models are ordered based on number of captions processed per video. Models with video-caption pretraining or utilizing significantly more captions than 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fair comparison. Numbers in parentheses () indicate the maximum number of frames used. See Sec. A.2 in appendix for detailed results.
Metric Category VideoAgent VideoTree LVNet
Avg. Frames $\downarrow$ - 24.6 98.0 24.0
Acc.(%) $\uparrow$ Knowledge 52.2 60.7 63.0
Film & TV 42.5 52.5 45.0
Sports Comp. 42.7 48.6 48.0
Artistic Perf. 47.5 51.6 53.0
Life Record 44.7 49.5 45.0
Multilingual 36.6 40.0 53.0
[…]
Average 46.4 53.1 52.4
Table 3: Evaluating on Very Long Videos. Comparison of LVNet (ours) with VideoAgent and VideoTree on the long split of VideoMME. LVNet uses the fewest frames while achieving the highest accuracy in three out of six categories and ranking second in overall performance, slightly below the best score. In the table, bold indicates the best performance, while underlined represents the second-best performance.
4.1 Experimental Setup
Datasets: Given the training free nature of our framework, we do not utilize any video datasets for training. Datasets are used purely for evaluation. We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema (Mangalam et al., 2023), NExT-QA (Xiao et al., 2021), and IntentQA (Li et al., 2023a). In addition, to further highlight the strength of our approach on longer videos, we include results on VideoMME’s long splitFu et al. (2024). These datasets are public available and can be used freely for academic research. The first dataset, EgoSchema, consists of 5031 questions and each video lasts three-minute and have multiple choice question. The second dataset, NExT-QA, is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer. These questions are further classified as Causal (Cau.), Temporal (Tem.), and Descriptive (Des.) and we evaluate on its validation set containing 4996 questions over 570 videos. The third dataset, IntentQA, is based on NExT-QA videos corresponding to temporal and causal reasoning quetions. It consists of 16k multiple-choice questions which are classified as Why?, How? or Before/After (B./A.). The fourth dataset, VideoMME, consists of very long videos—some up to one hour long, with an average duration of 44 minutes, and provides 900 Q&A.
Model Choices & Hyperparameters: For the HKS module, we use the ResNet-18 (He et al., 2016a) for the TSC, CLIP-B/16 (Ranasinghe et al., 2023) for the CKD and GPT-4o for the FKD. We select ResNet-18 and CLIP-B/16 due to their smaller models sizes—0.01B and 0.12B, respectively–which are significantly lighter compared to GPT-4o, whose model size is expected to be on the scale of 100B-1T. This makes them well-suited for filtering dense frames efficiently. In line with previous state-of-the-art work Wang et al. (2024d); Zhang et al. (2023); Wang et al. (2023), we employ GPT API, especially GPT-4o, for both VLM and LLM. This choice is driven by its cost-effectiveness and lighter computational requirements compared to GPT-4. GPT-4o is used as the VLM for generating captions and as the LLM for answering questions in our framework. We run TSC and CKD on a single NVIDIA RTX A5000, which takes approximately two hours to process 500 questions. We use the default hyperparameters for each vision/language module, as we only perform inference, and set the LLM temperature to 0 to ensure reproducibility. Also, We use single run for our experiments.
4.2 Evaluation
Model Avg. Frame Captions (Rounded) $\downarrow$
6 7 8 10 11 12 16 32 62 70
VideoAgent 58.4 - 63.2 / 60.2 - 60.8 / 57.4 - - - - -
VideoTree - 61.0 - 61.6 62.2 - - 64.4 66.2 67.0
LVNet (Ours) - - 64.4 - - 68.2 67.8 - - -
Templating Order Acc.
Temporal 65.2
Confidence 67.6
Hybrid (both) 68.2
TSC CKD FKD Acc.
62.6
64.5
65.8
68.2
Table 4: Ablation study on EgoSchema (Mangalam et al., 2023): We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA.
Quantitative Results:
We evaluate LVNet on the EgoSchema, NExT-QA, and IntentQA dataset and present our results in Table 2. Models with video-caption pretraining are de-emphasized in grey to ensure fairness with image-level pertaining. Models utilizing significantly more captions than the 12 frames are downplayed in light green to consider caption efficiency. For EgoSchema, we achieve 61.1% on the fullest, the highest among the models utilizing approximately 12 captions. This result outperforms VideoAgent, the next best model using 8.4 captions, by +7%, is on par with VideoTree while using only 1/5 of the captions, and outperforms TraveLER by +7.8% while utilizing only 12% of the captions.
We next evaluate on the NExT-QA dataset. This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs. Our approach achives state-of-the-art performance on this benchmark outperforming prior work among the models utilizing approximately 12 captions. In fact, our LVNet outperforms VideoAgent by +1.6%.
In the IntentQA dataset. LVNet outperforms all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames. In fact, LVNet shows a substantial improvement of +4.8% over the next best model, VideoTree, while using only 13% of the captions (12 vs. 90).
Lastly, Table 4 presents the performance of LVNet on VideoMME’s long split, which consists of videos up to one hour long and compare it to other models using keyframe selection methods. Our method (LVNet) demonstrates strong performance while utilizing only 24 frames, significantly fewer than VideoTree’s 98 frames. LVNet outperforms VideoAgent by +6.0% overall and achieves the highest accuracy in three out of six categories: Knowledge, Artistic Performance, and Multilingual. While VideoTree maintains a slight overall lead, LVNet’s ability to achieve comparable accuracy while processing only one-quarter of the frames highlights its efficiency in handling very long videos. To ensure a fair comparison, all models utilize GPT-4o.
Given the generative nature of VQA tasks as well as the limited availability and noisy nature of fully-annotated video VQA corpora, building generalizable fully-supervised models are challenging for these tasks. Nevertheless, we highlight how our zero-shot and video level training-free framework is competitive with the best supervised approaches on this dataset. This indicates the promise of utilizing pretrained models, especially those equipped with extensive world knowledge and reasoning skills from alternate modality specific learning (i.e. in our cases image domain VLMs and language domain LLMs).
Figure 4: Open-ended Responses from LVNet vs Uniform Sampling: The frames chosen by LVNet and the naive uniform sampling method are indicated with blue and red checkmarks, respectively. LVNet identifies both welding torches and measuring tapes, choosing the correct option, whereas uniform sampling only detects welding tools and selects the incorrect answer. The blue, red, and purple highlights correspond to the three main activities in the video—welding a handle, using a hammer, and using a measuring tape, respectively.
Qualitative Analysis of Hierarhical Keyframe Selector:
We compare the open-ended responses of LVNet and the uniform sampling method in Figure 4 to understand the effectiveness of the hierarchical keyframe selector in LVNet. The frames chosen by LVNet and the naive uniform sampling method are indicated by blue and red checkmarks in the images, respectively. LVNet selects frames at 5, 69, and 135 seconds by executing the hierarchical keyframe selector and generates captions based on those frames. When we feed the concatenated captions to the LLM to answer the given question: "Based on the video, what are the three main types of tools that C uses…" in an open-ended manner, the output identifies two main activities: welding torches and measuring tapes, among the three main activities described in Option 3 (welding handle, hammer, and measuring tape), which is the correct answer. This leads LVNet to choose the correct option.
In contrast, the uniform sampling method selects frames at 0, 16, and 32 seconds and generates captions based on those frames. Similarly, when we feed the concatenated captions to the LLM to answer the same question, the output identifies only one activity—welding tools—resulting in the selection of the incorrect option. This example highlights the importance of keyframe selection and demonstrates the effectiveness of hierarchical keyframe selection in LVNet.
4.3 Ablations
In this section, we present ablations on key design decisions such as the sorting order in FKD, the number of frames for captions, and the effect of different components in HKS. In all ablations, we use a subset of EgoSchema (Mangalam et al., 2023), composed of 500 videos. Additional ablations about Choice of LLM and Effect of Patch Size on Keyword Matching in CKD are in Appendix A.1
Visual Templating Order:
In visual templating, prioritizing frames by keyword confidence scores followed by reordering low-confidence frames based on timestamp proves more effective than using confidence scores or temporal order alone, as shown in Table LABEL:tab:ablation:visualtemplateorder. In this hybrid approach, high-confidence frames capture short but important segments of the video, while low-confidence keyframes, which are crucial but visually challenging for keyword matching, are temporally ordered to cover broader segments. This hybrid approach outperforms solely temporal ordering and solely confidence-based ordering by +3% and +0.6%, respectively.
Number of Frame Captions:
We performed an ablation study on the number of frame captions, comparing our approach to VideoAgent (Wang et al., 2024b) and VideoTree (Wang et al., 2024e) under similar low caption settings. As shown in Table LABEL:tab:ablation:numberofframecaptions, LVNet achieves the highest accuracy of 68.2% with 12 captions, outperforming VideoAgent (8.4 frames) and VideoTree (12 frames) by +8% and $\sim$+5.7%, respectively. We compare LVNet with VideoAgent+GPT-4o (8.1 frames) and VideoTree+GPT-4o (69.5 frames, $\times$5.8 more), both using GPT-4o for a fair comparison. We take GPT-4o numbers for VideoAgent and VideoTree from Yang et al. (2024) (more details in appendix A.2) and LVNet outperforms them by +5% and +1.2%, respectively.
Effect of Hierarchical Keyframe Modules:
Table LABEL:tab:ablation:effectofcomponents demonstrates the impact of incrementally adding the temporal scene clustering (TSC), coarse keyframe detector (CKD), and fine keyframe detector (FKD) modules. Without any of these modules, the model relies on uniform sampling and achieves 62.6%. When TSC is added and 12 frames are selected uniformly, the accuracy increases to 64.5%. Adding both TSC and CKD raises the accuracy to 65.8%. Finally, incorporating all three modules—TSC, CKD, and FKD—into the model, which is LVNet, results in an accuracy of 68.2%. This demonstrates the importance of including all modules in LVNet for optimal performance.
5 Conclusion
We proposed a novel approach for Long-form Video Question Answering (LVQA) that achieves state-of-the-art performance compared to the model using the similar-scale captions across 3 benchmarks datasets. Our Hierarchical Keyframe Selector demonstrates the effectiveness of keyframe selection in understanding a very long-form video QA. Additionally, we highlight the zero-shot capability for long-form video comprehension of our LVNet framework, which requires no video-level training. Our experiments showcase its significant advantage over previous methods.
Limitations
Despite the effectiveness of LVNet, as demonstrated by benchmark experiments and comprehensive ablations, our study has certain limitations, which we discuss below.
  • First, we acknowledge that we are unable to evaluate LVNet and other models with all available VLMs or LLMs due to computational constraints and high costs. However, we carefully select GPT-4o, a state-of-the-art LLM, for our main experiments and provide ablation studies comparing various LLMs (e.g. GPT-3.5, GPT-4, and GPT-4o) to other models to ensure a fair performance comparison, as presented in LABEL:tab:ablation:numberofframecaptions and LABEL:tab:ablation:choiceofllm.
  • Our hierarchical keyframe selector consists of three components: TSC, CKD, and FKD. While we demonstrated the effectiveness of each component in LABEL:tab:ablation:effectofcomponents, we did not have the time or resources to develop a unified module that could replace all three. Although this is beyond the scope of this paper, exploring a more efficient implementation that integrates these three modules into a single model would be an interesting direction for future research.
  • Like any LLM-based approach, LVNet is sensitive to prompting. To ensure the transparency, we provide examples of these prompts in Figure 4 and Figure A.7. We also plan to release the code to enable further exploration by other researchers.
  • Finally, we acknowledge that, as our approach is zero-shot, any inherent limitations or biases in the pretrained models may persist in the outputs of LVNet.
Acknowledgements:
This work was supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government [24ZB1200, Research of Human-centered autonomous intelligence system original technology]. This work was also supported by the Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No. RS-2024-00336738, Development of Complex Task Planning Technologies for Autonomous Agents, 100%)
Appendix
Figure A.5: Comparison of Keyframe Selection: Comparison of LVNet and VideoAgent in keyframe selection for video question answering. LVNet refines frames through a multi-stage process (TSC, CKD, FKD) to form a non-uniform keyframe distribution, capturing relevant moments tied to the query. In contrast, VideoAgent relies on uniform sampling and LLM-based frame selection, which fails to focus on crucial keyframes, leading to incorrect predictions. The final keyframe distributions illustrate LVNet’s ability to retrieve meaningful frames directly related to the answer, while VideoAgent selects irrelevant frames.
Appendix A.1 Additional Ablations
In this section, we present additional experiments conducted to inform the LVNet’s design. We have tested different LLMs and experimented with various scales of the visual feature map.
LLM Acc. (%)
GPT-3.5 61.0
GPT-4 65.4
GPT-4o 68.2
Patch Size Acc. (%)
1x1 63.6
7x7 66.2
14x14 68.2
Table A.5: Additional ablations experiments on EgoSchema (Mangalam et al., 2023): We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA. Default setting is highlighted.
Choice of LLM:
Table LABEL:tab:ablation:choiceofllm shows that GPT-4o outperforms both GPT-4 and GPT-3.5 by +2.8% and +7.2%, respectively. Given that GPT-4o is more cost-effective and lightweight compared to GPT-4, we have selected it as our default LLM.
Effect of Patch Size on Keyword Matching in CKD:
Table LABEL:tab:ablation:patchsizeinckd shows the effect of the scales of the patch sizes in the CKD. Since keywords can represent activities spanning the entire image or confined to a small region, we adjust the resolution of the visual feature map output from the spatially aware contrastive image pre-training (CLIP) network (Ranasinghe et al., 2023) to match keywords. Our findings show that higher resolutions lead to better accuracy. In LVNet, we use a 14$\times$14 feature map and determine the confidence level of the keyword by selecting the maximum value between the 14$\times$14 patches and the keyword’s text embedding.
Table A.6: Exact (Unrounded) Frame Caption Counts. These values supplement the rounded numbers in Table LABEL:tab:ablation:numberofframecaptions All models are based on either GPT-4o or GPT-4.
Model Avg. Frame Captions $\downarrow$
6.4 7.1 8 8.1 8.4 9.7 10.7 11 11.3 12 16 32 62.4 69.5
VideoAgent 58.4 - - 63.2 60.2 - 60.8 57.4 - - - - - -
VideoTree - 61.0 - - - 61.6 - - 62.2 62.5 66.2 64.4 66.2 67.0
LVNet (Ours) - - 64.4 - - - - - - 68.2 67.8 - - -
Exact Frame Caption Counts:
For completeness, Table A.6 lists the original (unrounded) frame caption counts and corresponding accuracies for VideoAgentWang et al. (2024b), VideoTreeWang et al. (2024e), and LVNet (ours). These values supplement the rounded numbers presented in Table LABEL:tab:ablation:numberofframecaptions of the main text.
Appendix A.2 Extended results on NExT-QA and IntentQA
We present extended zero-shot evaluation results on NExT-QA in Table A.7, comparing LVNet with prior zero-shot models across different task categories: causal, temporal, and descriptive reasoning. Models are ordered based on the number of captions processed per video, highlighting the trade-offs between caption efficiency and performance.
LVNet achieves state-of-the-art performance with an overall accuracy of 72.9%, outperforming most models while using only 12 captions per video. Notably, it attains 75.0% on causal reasoning, which is the highest among all models evaluated. For temporal reasoning, LVNet achieves 65.5%, remaining competitive despite using significantly fewer captions than models like VideoTree (56 captions) and LangRepo (90 captions). In descriptive reasoning, LVNet reaches 81.5%, matching VideoTree while processing significantly fewer captions.
Compared to VideoAgent, the closest competing model in terms of caption efficiency (8.4 captions), LVNet demonstrates a substantial performance gain across all categories, with a +2.8% improvement in overall accuracy. While models like VideoTree and TraveLER show strong performance, they process significantly more captions (56 and 65, respectively), indicating that LVNet achieves a superior balance between efficiency and accuracy.
We present extended zero-shot evaluation results on IntentQA in Table A.8, comparing LVNet with prior zero-shot models across different reasoning categories: Why?, How?, and B.A. (Before/After). Models are ordered based on the number of captions processed per video, highlighting the balance between caption efficiency and performance.
LVNet achieves an overall accuracy of 71.7%, outperforming all models while using only 12 captions per video. It achieves 75.0% on the Why? category, 74.4% on the How? category, and 62.1% on the B.A. category. Compared to VideoTree, which processes 56 captions and achieves an overall accuracy of 66.9%, LVNet outperforms it by +4.8% while using significantly fewer captions. Similarly, LangRepo and LLoVi, which process 90 captions, achieve overall scores of 59.1% and 64.0%, respectively, further demonstrating LVNet’s caption efficiency.
To ensure fairness, models that utilize video-caption pretraining or process substantially more captions than LVNet are de-emphasized in grey or downplayed in light green in Table A.7 and Table A.8. We adopt the reported GPT-4o results for VideoAgent and VideoTree in Section 4 and LABEL:tab:ablation:numberofframecaptions from the VCA Yang et al. (2024), but do not compare directly against VCA for two reasons: (1) the VCA paper does not provide code or implementation details (e.g., inference speed), making replication infeasible; and (2) its reported results cover only a subset of Egoschema, preventing a fair comparison to our approach on the full-scale EgoSchema, NExT-QA, and IntentQA. Overeall, these clarifications, alongside the results in Table A.7 and Table A.8, underscore LVNet’s effectiveness in achieving high accuracy while maintaining computational efficiency.
Model Cap. Cau. (%) Tem. (%) Des. (%) All (%)
IG-VLM (Kim et al., 2024) - 69.8 63.6 74.7 68.6
Tarsier (Wang et al., 2024a) - - - - 79.2
VideoAgent (Wang et al., 2024b) 8.2 72.7 64.5 81.1 71.3
MVU (Ranasinghe et al., 2024) 16 55.4 48.1 64.1 55.2
MoReVQA (Min et al., 2024) 16 70.2 64.6 - 69.2
VFC (Momeni et al., 2023) 32 45.4 51.6 64.1 51.5
SeViLA (Yu et al., 2024) 32 61.3 61.5 75.6 63.6
VideoTree (Wang et al., 2024e) (56) 75.2 67.0 81.3 73.5
ProViQ (Choudhury et al., 2023) 60 - - - 64.6
TraveLER (Shang et al., 2024) (65) 70.0 60.5 78.2 68.2
LangRepo (Kahatapitiya et al., 2024) 90 64.4 51.4 69.1 60.9
LLoVi (Zhang et al., 2023) 90 69.5 61.0 75.6 67.7
LVNet (ours) 12 75.0 65.5 81.5 72.9
Table A.7: Extended results on NExT-QA (Xiao et al., 2021). We compare LVNet against prior zero-shot models across different reasoning categories: causal, temporal, and descriptive. LVNet achieves an overall accuracy of 72.9% while using only 12 captions per video, demonstrating strong performance across all reasoning types. Notably, it outperforms all models in causal reasoning (75.0%) and matches the best performance in descriptive reasoning (81.5%), despite processing significantly fewer captions than models like VideoTree (56 captions) and TraveLER (65 captions). Models that utilize video-caption pretraining or process substantially more captions than LVNet are de-emphasized in gray or downplayed in light green to ensure fairness in comparison. Numbers in parentheses () indicate the maximum number of frames used.
Model Cap. Why? (%) How? (%) B./A. (%) All (%)
IG-VLM (Kim et al., 2024) - - - - 65.3
SeViLA (Yu et al., 2024) 32 - - - 60.9
VideoTree (Wang et al., 2024e) (56) - - - 66.9
LangRepo (Kahatapitiya et al., 2024) 90 62.8 62.4 47.8 59.1
LLoVi (Zhang et al., 2023) 90 68.4 67.4 51.1 64.0
LVNet (ours) 12 75.0 74.4 62.1 71.7
Table A.8: Extended results on IntentQA (Li et al., 2023a). We compare LVNet against prior zero-shot models across different reasoning categories: Why?, How?, and B.A. (Belief/Action). LVNet achieves an overall accuracy of 71.7%, surpassing all models while using only 12 captions per video. It reaches 75.0% in the Why? category, 74.4% in the How? category, and 62.1% in the B.A. category. Compared to VideoTree, which processes 56 captions and achieves 66.9% accuracy, LVNet outperforms it by +4.8% while using significantly fewer captions. Additionally, LVNet demonstrates superior reasoning-based performance compared to LangRepo (90 captions, 59.1%) and LLoVi (90 captions, 64.0%). Models with video-caption pretraining or utilizing significantly more captions than 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fairness with image-level pretraining or highlight caption efficiency. Numbers in parentheses () indicate the maximum number of frames used.
Appendix A.3 Algorithms in Detail
Our algorithms are presented in full detail in Algorithm 1, Algorithm 2, and Algorithm 3. TSC in Algorithm 1 extracts per-frame visual features using ResNet-18, followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets. Within each of the $n$ sets, we uniformly sample $\leq\tau$ frames, obtaining a total of $T_{a}\leq\tau\times n$ frames. For example, LVNet sets $\psi=5,\lambda=12,\tau=18$, resulting in approximately $n\sim 25$ and $T_{a}\sim 390$ on the EgoSchema dataset. CKD in Algorithm 2 selects top $L$ frames based on similarity/confidence scores, which are calculated using cosine similarity between frames and keywords with CLIP-B/16. LVNet employs $L=32,len(K)\leq 25$ on the EgoSchema dataset. FKD in Algorithm 3 sorts frames and their corresponding keywords by confidence scores, and reorder the $K$ frames with the lowest scores temporally. It groups frames sequentially into visual templates, each consisting of $N$ frames. From each template, the $M$ frames and keywords most relevant among the $N$ pairs are selected using GPT-4o. We set $L=32,K=16,N=8,M=3$.
1
21:Require: ResNet-18 He et al. (2016b) pretrained on imagenet dataset $f$, frame list $\textbf{List}_{frame}$, image index list $\textbf{List}_{index}\in\{1,\dots,N\}$, minimum number of list length $\psi$, temperature $\lambda$, number of sample $\tau$, function to find index of x in list w index(x, w), and function to sort list sort(List)
3
2:for all $img^{i}$ in $\textbf{List}_{frame}$ do
43:   $\mathbf{F}^{i}\leftarrow f(img^{i})$
4:   $\textbf{List}_{feat}$.insert($\mathbf{F}^{i}$)
55:end for
6:for all $\mathbf{F}^{i}$ in $\textbf{List}_{feat}$ do
67:   $\textbf{List}_{dist}\leftarrow\frac{\sum_{y}\sum_{x}{\sqrt{(\mathbf{F_{i}}-\textbf{List}_{feat})^{2}}}}{x\times y}$
8:   $\textbf{M}_{dist}$.insert($\textbf{List}_{dist}$)
79:end for
10:while length of $\textbf{List}_{index}>\psi$ do
11:   $\textbf{List}_{sample}\leftarrow\emptyset$
12:   $\textbf{List}_{\delta}\leftarrow\emptyset$
813:   $i\leftarrow\textbf{List}_{index}$.pop(00)
914:   $\textbf{p}^{i}\leftarrow\texttt{softmax}(\textbf{M}^{i}_{dist})$
1015:   $\mu_{\mathbf{\textbf{p}^{i}}},\sigma_{\mathbf{\textbf{p}^{i}}}\leftarrow\texttt{mean}(\textbf{p}^{i}),\texttt{std}(\textbf{p}^{i})$
1116:   $\beta\leftarrow\mu_{\mathbf{\textbf{p}^{i}}}-\sigma_{\mathbf{\textbf{p}^{i}}}\sum_{i=0}{e^{1-i/\lambda}}$
17:   for all prob in $\textbf{p}^{i}$ do
18:      if prob < $\beta$ then
19:         $\textbf{List}_{selected}$.insert($\textbf{index}(\texttt{prob},\textbf{p}^{i})$)
20:      end if
1221:   end for
22:   for all $\gamma$ in $\textbf{List}_{selected}$ do
1323:      $\delta$leftarrow $\gamma~{}th$ value in $\textbf{List}_{index}$
24:      $\textbf{List}_{\delta}$.insert($\delta$)
25:      $\textbf{List}_{index}$.pop($\gamma$)
1426:   end for
1527:   $\textbf{List}_{\delta}$.insert($i$)
1628:   $\textbf{List}_{sample}$leftarrow sample $\tau$ items from $\textbf{List}_{\delta}$
1729:   $\textbf{sort}(\textbf{List}_{sample})$
30:   for all $frame^{j}$ in $\textbf{List}_{frame}$ do
31:      if $j$ in $\textbf{List}_{sample}$ then
32:         Outputs.insert($frame^{j}$)
33:      end if
1834:   end for
35:end while
Algorithm 1 Temporal Scene Clustering
1
21:Require: keyword set $\mathbf{K}$, image set $\mathbf{I}$, total length of selected image set $L$, function to calculate similarity matrix $\textbf{sim}(\mathbf{K},\mathbf{I})$, function to sort similarity matrix and return indices $\textbf{sort}(\mathbf{S})$
3
42:$\mathbf{S}\leftarrow\textbf{sim}(\mathbf{K},\mathbf{I})$
53:$\mathbf{S}_{sorted},\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
4:Initialize $\mathbf{P}_{best}$ as an empty list
65:Initialize $\mathbf{I}_{selected}$ as an empty set
6:while length of $\mathbf{I}_{selected}<L$ do
7:   for $k\in\mathbf{K}$ do
8:      for $i\in\mathbf{I}$ do
9:         $i_{index}\leftarrow\mathbf{idx}_{sorted}[k][i]$
10:         if $i_{index}$ not in $\mathbf{I}_{selected}$ then
11:            $\mathbf{P}_{best}$.insert($k,i_{index}$)
12:            $\mathbf{I}_{selected}$.insert($i_{index}$)
13:            break
14:         end if
15:      end for
16:      if length of $\mathbf{I}_{selected}\geq L$ then
17:         break
18:      end if
19:   end for
720:end while
821:return $\mathbf{P}_{best}$
Algorithm 2 Keyword-Image Matching Process in CKD
1
[…]
3
42:$\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
[…]
64:$\mathbf{idx}_{temporal}\leftarrow\textbf{temporal\_order}(\mathbf{idx}_{low\_sim})$
[…]
86:$\mathbf{I}_{ordered},\mathbf{K}_{ordered}\leftarrow\mathbf{I}[\mathbf{idx}_{final}],\mathbf{K}[\mathbf{idx}_{final}]$
[…]
8:for each $\mathbf{set}\in\mathbf{sets}$ do
9:   $\mathbf{I}_{selected}\leftarrow\texttt{select\_top\_M}(\mathbf{set},M)$
1010:end for
$1$:return $\mathbf{I}_{selected}$
Algorithm 3 Fine Keyframe Detection Process (FKD)
Appendix A.4 Prompting: Fine Keyframe Detector
We prompt the VLM to select frames that are most compatible with the list of given keywords. Each template image contains 8 images, and their order is described in language (e.g. top left to right, bottom left to right) and the VLM outputs the selected images according to our prompting as described in Figure A.7.
Figure A.7: Prompt for Fine Keyframe Detection: The figure illustrates the input image, the prompt provided to the VLM, and the output. The input image represents a visual template composed of eight frames, and the prompt requests the three best frames along with their corresponding keywords. The output displays the top three selected frames and their associated keywords.
Appendix A.5 Comparison with Other Keyframe Selection Methods
We aim to highlight the main advantage of the Hierarchical Keyframe Selector over other existing keyframe selection methods. Models like VideoAgent, VideoTree, and TraveLER provide useful comparisons, as they utilize keyframe selection mechanism with similar or different scale of frames. VideoAgent and TraveLER rely on uniform frame selection in the first iteration without analyzing the entire video even though they perform non-uniform sampling in the next iterations. They identify important segments based solely on these initial frames and the LLM’s response, which can be problematic if the initial uniformly selected frames are not representative of the entire video or if the LLM misinterprets the captions and prompts. In such cases, the LLM might incorrectly identify segments for further analysis. If the LLM fails to pinpoint the correct segment initially, the entire process can break down because subsequent frames will be similar to the first set, leading the LLM to continuously select frames within or near the initial segment. Additionally, for videos that are as challenging or more difficult than EgoSchema in terms of temporal complexity and activities, existing keyframe selection models such as VideoAgent, VideoTree, and TraveLER may require numerous iterations by running heavy visual/language models to finalize keyframes selection. This results in higher computational and latency costs, as it necessitates numerous runs of resource-intensive VLM and LLM models.
In contrast, our method analyzes the entire video with high frame rates using a lightweight ResNet-18 (He et al., 2016a) and segments the video non-uniformly based on scene continuity. We then select several frames in each segment by measuring feature similarity between frame features and keywords using the CLIP-B/16 (0.12B) (Ranasinghe et al., 2023) which is lighted than VideoAgent’s EVA-CLIP-8Bplus (8B). By reviewing the entire video and non-uniformly selecting keyframes based on scene continuity and similarity scores, these keyframes accurately represent the question-based important frames distribution in the entire video. Furthermore, we use VLM for a fine-grained selection of keyframes, improving keyframe selection when CLIP-B/16 struggles to understand detailed atomic activities in the frames. By hierarchically segmenting the video with different modules, the resulting segments and keyframes are more reliable than those from VideoAgent. Even with more challenging videos, our process only needs to go through the video once to collect keyframes, maintaining computational efficiency.
Figure A.5 visualizes the differences of the keyframe selection mechanism bewtween LVNet and VideoAgent. On the left, LVNet begins with uniformly sampled frames and filters them through multiple stages, resulting in a non-uniform distribution of frames over time. First, the temporal scene clustering (TSC) selects somes frames that represent temporally distinct activities. Next, the coarse keyframe detector (CKD) targets frames most relevant to the question. Finally, the fine keyframe detector (FKD) further refines this selection to ensure the keyframes accurately capture the activity in question. As a result, LVNet produces 12 frames, with 8 of them (67%) directly depicting "usage of phones," which is the correct answer and leads the model to select the right option. On the right, VideoAgent also starts with the uniform frames but relies on a LLM to request additional frames. Since the initial frames do not capture enough relevant content, the LLM again selects frames uniformly, adding more irrelevant samples that lack the crucial information about "usage of phones." As a result, VideoAgent ultimately selects the wrong option.